### v1

# Divergence-Kernel method for linear responses and diffusion models

Abstract.
We derive the divergence-kernel formula for the linear response (parameter-derivative of marginal or stationary distributions) of random dynamical systems, and formally pass to the continuous-time limit. Our formula works for multiplicative and parameterized noise over any period of time; it does not require hyperbolicity. Then we derive a pathwise Monte-Carlo algorithm for linear responses. With this, we propose a forward-only diffusion generative model and test on simple problems.
AMS subject classification numbers. 60H07, 60J60, 65D25, 65C30, 65C05.
Keywords. Linear response, Divergence method, Likelihood ratio, SDE, Diffusion model.
1. Introduction
1.1. Main results
This paper rigorously derives the divergence-kernel formula for the linear response of discrete-time random dynamical systems. Then we formally pass to the continuous-time limit. Our result gives a fundamental tool for the optimization of diffusion processes; in particular, from now on, we can use the gradient-based method to optimize marginal or stationary densities with respect to the diffusion coefficients. In particular, we propose a simple framework for diffusion generative models. Our main theoretical result is
{theorem}
[formal divergence-kernel formula for linear responses of SDEs] Consider the SDE in $\mathbb{R}^{M}$ parameterized by $\gamma$,
$dx^{\gamma}_{t}=F^{\gamma}_{t}(x^{\gamma}_{t})dt+\sigma^{\gamma}_{t}(x^{\gamma}_{t})dB_{t},$
For any backward adapted process $\{\alpha_{t}\}_{t=0}^{T}$, let $\{\nu_{t}\}_{t=0}^{T}$ be the forward covector process at $\gamma=0$ such that
$\begin{split}d\nu=\left((\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha)\nu-\nabla\textnormal{div}F+\nabla^{2}\sigma\nabla\sigma+\nabla\sigma\Delta\sigma\right)dt-(\nabla\sigma\nu^{T}+\nabla^{2}\sigma+\alpha\sigma^{-1})dB.\end{split}$
with initial condition $\nu_{0}=\nabla\log h_{0}(x_{0})$, where $h_{t}$ is the probability density of $x_{t}$. Then $\nabla\log h_{t}(x_{t})=\mathbb{E}\left[\nu_{t}\middle|x_{t}\right]$, and
$\begin{split}\delta\log h^{\gamma}_{T}(x_{T})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}+\int_{t=0}^{T}\delta\sigma^{\gamma}\nabla\sigma^{T}\nu-\delta F^{\gamma T}\nu-\textnormal{div}\delta F^{\gamma}+\delta\sigma^{\gamma}\Delta\sigma+\nabla\sigma^{T}\nabla\delta\sigma^{\gamma}dt\right.\\ \left.-\int_{t=0}^{T}(\delta\sigma^{\gamma}\nu^{T}+\nabla\delta\sigma^{\gamma T})dB\middle|x_{T}\right].\end{split}$
Here $\Delta\sigma$ is the Laplacian of $\sigma$, $B$ is the Brownian motion, and the SDE is Ito. The time integrations are the limits of Equations˜9 and 10. More generally, $\nu$ is an estimator for the score of the SDE at $\gamma=0$, which can be replaced by other divergence-kernel expressions from [24].
The main advantages of this result are:
  1. (1)
    It computes derivative of the marginal density, not an averaged observable;
  2. (2)
    The expression in pathwise, so we can compute by Monte-Carlo type algorithms in high-dimensions;
  3. (3)
    The diffusion coefficient $\sigma$ can depend on the parameter $\gamma$ and the location $x$;
  4. (4)
    The size of the estimator does not grow exponentially to $T$, so we can easily extend our result to stationary distributions (see ˜Remark);
  5. (5)
    It does not require hyperbolicity (in the sense of dynamical systems).
As a less obvious application, we use this result to construct a new diffusion generative model, where we directly optimize the marginal distribution to reduce its KL-divergence with respect to the data. Importantly, we can optimize the diffusion coefficients of multiplicative noise.
1.2. Literature review
The marginal or stationary distribution of a random dynamical system is of central interest in applied sciences. We are interested in the parameter-derivative $\delta\log h_{t}$, where $h_{t}$ is the probability density at time $t$. There are three basic methods for expressing and computing the linear response of random dynamical systems: the path-perturbation method (shortened as the path method), the divergence method, and the kernel-differentiation method (shortened as the kernel method). The relation and difference among the three basic methods can be illustrated in a one-step system, which is explained in [25].
The path-perturbation method computes the path-perturbation under fixed randomness (such as Brownian motion), then averages over many realizations of randomness. This is also known as the ensemble method or the stochastic gradient method [7, 15]. It also includes the backpropagation method, which is the basic algorithm for machine learning. However, it is expensive for chaotic or unstable system; the work-around is to artificially reduce the size of the path-perturbation, such as shadowing or clipping methods [26, 29, 30, 8], but they all introduce systematic errors. Moreover, it computes the derivative of an averaged observable, but does not give the derivative of the probability density.
The divergence method is also known as the transfer operator method, since the perturbation of the measure transfer operator is some divergence [10, 36, 38]. However, it is expensive for high-dimensional systems with contracting directions.
The kernel-differentiation method works only for random systems. It is also known as the Cameron-Martin-Girsanov theorem [4, 17] or the likelihood ratio method [33, 31, 11]. However, it cannot handle multiplicative noise or perturbation on the diffusion coefficients. It is also expensive when the noise is small. We gave an ergodic version and a foliated version in [25].
Mixing two basic methods can overcome some of their major shortcomings. Our path-divergence method (also known as the fast response formula) gives the pointwise expression of linear responses of hyperbolic deterministic chaos [21, 27, 19, 26, 20, 28]. It is good at high-dimensional and low-noise to deterministic systems. It can be used to solve the optimal response problem in the hyperbolic setting [9]. However, it does not work when the hyperbolicity is poor [1, 37], and only the unstable part involves derivatives of densities. The path-divergence idea was also used in the proof of the basic path-perturbation or divergence linear response formulas for hyperbolic systems [5, 34, 13, 12, 2].
Our path-kernel method in [23] gives the pathwise linear response formula of the diffusion coefficients. It is good at systems with not too small noise and not too much unstableness, it does not require hyperbolicity. However, it can be expensive when the noise is small and unstableness is big. Moreover, only part of the expression is derivative on the terminal densities. The path-kernel idea was also used in the Bismut-Elworthy-Li formula for the derivative with respect to the initial conditions [3, 6, 32].
Our divergence-kernel method in [24] gives the score (derivative of marginal density) of random dynamical systems and SDEs. It works for non-hyperbolic and multiplicative noise systems. It is good at systems with not too much contraction and not too little noise. However, it can be expensive when the noise is small and the contraction is strong. We use this result for the forward-backward diffusion models via nonlinear SDEs [18]. This paper extends this result to linear responses.
Our new linear response formula opens up the possibilities of redesigning the current framework of diffusion models. The current structure of the diffusion model involves a forward process and a backward process [35, 16]. Part of the reason is that previously we did not have a gradient tool to optimize marginal distributions with respect to the diffusion coefficient or under multiplicative noise. If we can directly compute such a gradient, we can directly optimize the diffusion process so that the marginal distribution matches the data. We will demonstrate this idea on some simple problems.
1.3. Structure of paper
Section˜2 defines some basic notation and states some basic assumptions on regularities. Section˜3.1 derives our the divergence formula for the linear response of one-step random systems. Section˜3.2 derives our results for discrete-time random dynamical systems by combining the one-step divergence linear response formula with the multi-step divergence-kernel formula for scores. Section˜3.3 recalls some useful facts from [24] about scores for SDEs. Section˜3.4 formally passes the linear response formula to the limit of SDEs. Section˜3.5 discusses about how to use this result.
Section˜4 uses the new formulas to numerically compute the linear responses of two model systems. Section˜4.1 considers a 1-dimensional process with multiplicative noise. For this 1-d example, we can plot the parameter-derivative of the density. Section˜4.2 computes the linear response of the stationary measure of the 40-dimensional Lorenz 96 model with multiplicative noise. These examples can not be solved by previous methods since the system is high-dimensional, has multiplicative noise, has contracting directions, and is non-hyperbolic. Moreover, compared with our path-kernel methods, here we can compute the derivative of the density.
Section˜5 proposes a diffusion model which involves only two forward processes. It uses the divergence-kernel formula (not the backpropagation method, which belongs to the path-perturbation method) to perform a gradient descent on the parameters of an SDE. Its loss function is the KL-divergence between the data and the marginal distribution of the SDE. The logic is more straightforward than the current diffusion models, and the memory cost is less.
Appendix˜A briefly goes through the kernel-differentiation method and explains its problem when applied to linear responses of high-dimensional systems with respect to scalar diffusion coefficients. We also derive the conventional likelihood ratio formula for additive noise systems using our notation.
2. Notations and assumptions
We define some geometric notations for spatial derivatives. In this paper, the ambient space is the $M$-dimensional Euclidean space $\mathbb{R}^{M}$. We denote both vectors and covectors by column vectors in $\mathbb{R}^{M}$. The product between a covector $\nu$ and a vector $v$ is denoted by $\cdot$, that is,
$\begin{split}\nu\cdot v:=v\cdot\nu:=\nu^{T}v:=v^{T}\nu.\end{split}$
Here $v^{T}$ is the transpose of matrices or (co)vectors. Then we use $\left\langle,\right\rangle$ to denote the inner-product between two vectors
$\begin{split}\left\langle u,v\right\rangle:=u^{T}v.\end{split}$
Note that $\Delta B$} may be either a vector or a covector.
The score function is $\nabla\log{h_{T}}=\frac{\nabla h_{T}}{h_{T}}$, where
$\nabla(\cdot):=\frac{\partial(\cdot)}{\partial x},\quad\nabla_{v}(\cdot):=\nabla(\cdot)v:=\frac{\partial(\cdot)}{\partial x}v,$
Here $\nabla_{Y}X$ denotes the (Riemann) derivative of the tensor field $X$ along the direction of $Y$. It is convenient to think that $\nabla$ always adds a covariant component to the tensor. The divergence of a vector field $X$ is defined by
$\begin{split}\textnormal{div}X:=\sum_{i=1}^{M}\nabla_{e_{i}}X,\end{split}$
where $e_{i}$ is the $i$-th unit vector of the canonical basis of $\mathbb{R}^{M}$.
For a map $g$, let $g_{*}$ be the Jacobian matrix, or the pushforward operator on vectors; readers may also denote this by $\nabla g$. We define the covector $\textnormal{div}g_{*}$, the divergence of the Jacobian matrix $g_{*}$, as
(1) $\begin{split}\textnormal{div}g_{*}:=\nabla\log\left|g_{*}\right|:=\frac{\nabla\left|g_{*}\right|}{\left|g_{*}\right|},\end{split}$
where $\left|g_{*}\right|$ is the determinant of the matrix.
For discrete-time results in Section˜3.2, we assume that
Assumption.
The density of the initial distribution $h_{0}$, the probability kernel $k$, drift $f$, and diffusion $\sigma$ are $C^{1}$ jointly in $\gamma$ and $x$.
Our derivation is rigorous for the discrete-time case, but is formal when passing to the continuous-time limit, so we do not list the technical assumptions here, and leave the rigorous proof to a later paper. We just assume that all integrations, averages, and change of limits are legit. We still organize our continuous-time results by lemmas and theorems, but this is just a style to organize the paper and does not imply rigorousness.
3. Derivation of the divergence-kernel formula for linear responses
There are two basic ingredients to our work. The first is the divergence formula for $\delta\log h^{\gamma}_{1}$ in the one-step system, which we shall derive now. The second is the divergence-kernel formula for the scores in many-step systems and its continuous-time limit; this result is from [24].
3.1. Divergence formula One-step system
Let $x^{\gamma}_{0}\sim h^{\gamma}_{0}$, $b^{\gamma}_{0}\sim k$, the one-step dynamics is
$x^{\gamma}_{1}=f^{\gamma}_{0}(x^{\gamma}_{0})+\sigma^{\gamma}_{0}(x^{\gamma}_{0})b_{0}.$
Here $\gamma$ is the parameter; when we omit $\gamma$ in our notation, it takes the default value 0. We sometimes omit the subscripts that indicate the time step of $f$ and $\sigma$, which are the same as the time step of the variable $x$. Let $h^{\gamma}_{1}$ denote the density of $x^{\gamma}_{1}$, which has a natural expression:
(2) $h^{\gamma}_{1}(x_{1})=\int h^{\gamma}_{0}(x_{0})p^{\gamma}_{0}(x_{0},x_{1})dx_{0}.$
where the probability kernel is
$\begin{split}p^{\gamma}_{0}(x_{0},x_{1}):=(\sigma^{\gamma}_{0})^{-M}(x_{0})k\left(\frac{x_{1}-f^{\gamma}_{0}(x_{0})}{\sigma^{\gamma}_{0}(x_{0})}\right).\end{split}$
Change the variable $x_{0}$ to $b_{0}=(x_{1}-f^{\gamma}_{0}(x_{0}))/\sigma^{\gamma}_{0}(x_{0})$ under fixed $x_{1}$. Define
$\begin{split}x_{1}=g^{\gamma}_{b_{0}}(x_{0}):=f^{\gamma}_{0}(x_{0})+\sigma^{\gamma}_{0}(x_{0})b_{0}\end{split}$
Since $x_{1}$ is fixed, $x_{0}$ and $b_{0}$ are related by
$\begin{split}0=dx_{1}=g^{\gamma}_{b_{0}*}(x_{0})dx_{0}+\sigma^{\gamma}_{0}(x_{0})db_{0},\end{split}$
where $g^{\gamma}_{b_{0}*}:=\nabla f^{\gamma}+b_{0}(\nabla\sigma^{\gamma})^{T}$ is the pushforward operator of $g_{b_{0}}$. Changing the variable causes a nontrivial Jacobian determinant.
$\begin{split}\left|\frac{db_{0}}{dx_{0}}\right|=(\sigma_{0}^{\gamma})^{-M}(x_{0})\left|g^{\gamma}_{b_{0}*}(x_{0})\right|,\end{split}$
where $\left|g^{\gamma}_{b_{0}*}\right|$ is the determinant. So we can express $h^{\gamma}_{1}$ by an integration over $b_{0}$,
(3) $\begin{split}h^{\gamma}_{1}(x_{1})=\int h^{\gamma}_{0}(x^{\gamma}_{0})k(b_{0})\left|g^{\gamma}_{b_{0}*}(x^{\gamma}_{0})\right|^{-1}db_{0},\quad\textnormal{where}\quad x^{\gamma}_{0}=(g^{\gamma}_{b_{0}})^{-1}(x_{1}).\end{split}$
For now, we assume $g^{\gamma}_{b_{0}}$ is injective, since this is the case when we discretize SDEs; if it is $n$-to-1, then we need to further sum over its preimage.
{lemma}
[Divergence formula for one-step linear response]
$\begin{split}\delta\log h^{\gamma}_{1}(x_{1})=\mathbb{E}\left[\frac{\delta h^{\gamma}_{0}+\nabla_{\delta x^{\gamma}_{0}}h_{0}}{h_{0}}(x_{0})-\frac{\delta\left|g^{\gamma}_{b_{0}*}\right|+\nabla_{\delta x^{\gamma}_{0}}\left|g_{b_{0}*}\right|}{\left|g_{b_{0}*}\right|}(x_{0})\middle|x_{1}\right],\end{split}$
where $\delta x^{\gamma}_{0}=-g^{-1}_{b_{0}*}\delta g^{\gamma}_{b_{0}}(x_{0}).$
Proof.
Differentiate Equation˜3 by $\gamma$, note that the dummy variable $b_{0}$ is not differentiated; the derivative hits $h^{\gamma}_{0}$, $g^{\gamma}_{b_{0}*}$, and $x^{\gamma}_{0}$. We get
$\begin{split}\delta h^{\gamma}_{1}(x_{1})=\int\left(\frac{\delta h^{\gamma}_{0}+\nabla_{\delta x^{\gamma}_{0}}h_{0}}{h_{0}}(x_{0})-\frac{\delta\left|g^{\gamma}_{b_{0}*}\right|+\nabla_{\delta x^{\gamma}_{0}}\left|g_{b_{0}*}\right|}{\left|g_{b_{0}*}\right|}(x_{0})\right)h_{0}(x_{0})k(b_{0})\left|g_{b_{0}*}(x_{0})\right|^{-1}db_{0},\end{split}$
To get $\delta x^{\gamma}_{0}$, differentiate $x_{1}=g^{\gamma}_{b_{0}}(x^{\gamma}_{0})$, note that $x_{1}$ is fixed, we get
$\begin{split}0=\delta g^{\gamma}_{b_{0}}(x_{0})+g_{b_{0}*}\delta x^{\gamma}_{0}.\end{split}$
Hence, $\delta x^{\gamma}_{0}=-g^{-1}_{b_{0}*}\delta g^{\gamma}_{b_{0}}(x_{0}).$ Then substitute to prove the lemma. ∎
We do not work on the divergence-kernel method for linear responses in one-step, since it is not directly used. The reason is that the kernel formula for linear responses (see Section˜A.1 in Appendix˜A) has a term with factor $M$, which can be expensive (since large terms require many samples to compute averages) for important applications. The kernel formula for scores does not have this large term. Hence, we choose to use the divergence formula for the linear response and the divergence-kernel formula for the scores.
That being said, we can certainly mix the kernel and divergence linear response formulas in one-step, then repeatedly apply to many-step systems, and then pass to the continuous-time limit. That approach might be beneficial if we later work with anisotropic noise, where $\sigma$ is a matrix, and each parameter controls only a low-dimensional structure, so we do not have the $M$ factor. Another scenario where this approach might be more useful is that the noise is low-dimensional. Bot situations incur other difficulties, so we will defer to later papers.
3.2. Discrete-time divergence-kernel formula for linear responses
Starting from the initial density $x^{\gamma}_{0}\sim h^{\gamma}_{0}$, consider the discrete-time random dynamical system
(4) $\begin{split}x^{\gamma}_{n+1}=f^{\gamma}_{n}(x^{\gamma}_{n})+\sigma^{\gamma}_{n}(x^{\gamma}_{n})b_{n},\end{split}$
where $b_{n}$ is i.i.d. distributed according to the kernel density $k$. Denote the density of $x^{\gamma}_{n}$ by $h^{\gamma}_{n}$. The system runs for a total of $N$ steps.
We use the divergence formula in one-step, then substitute the recent divergence-kernel formula for the score at $\gamma=0$. Note that the score formula contains the kernel-differentiation formula for the score, but not the kernel-differentiation formula for linear responses. Hence, we do not have the term involving $M$ which can potentially get very large for scalar diffusion coefficients. First, recall the following:
Definition 1.
We say that a process $\{\alpha_{n}\}_{n=0}^{N}$ is backward adapted, or adapted to the backward filtration, if $\alpha_{n}$ belongs to the $\sigma$-algebra $\sigma(x_{n},\ldots,x_{N})$.
{lemma}
[N-step forward divergence-kernel formula for score [24]] For any backward adapted process $\{\alpha_{n}\}_{n=0}^{N}$, let $\{\nu_{n}\}_{n=0}^{N}$ be the forward covector process
$\begin{split}\nu_{0}=\nabla\log h_{0}(x_{0}),\quad\textnormal{}\quad\nu_{n+1}=(1-\alpha_{n+1})g_{b_{n}}^{*-1}\left(\nu_{n}-\textnormal{div}g_{b_{n}*}(x_{n})\right)+\frac{\alpha_{n+1}\nabla\log k(b_{n})}{\sigma_{n}(x_{n})}.\end{split}$
Then for any $N\geq 0$,
$\begin{split}\nabla\log h_{N}(x_{N})=\mathbb{E}\left[\nu_{N}\middle|x_{N}\right].\end{split}$
{theorem}
[N-step divergence-kernel formula for linear responses]
$\begin{split}\delta\log h^{\gamma}_{N}(x_{N})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}+\sum_{n=0}^{N-1}\left(\nu_{n}-\nabla\log|g_{b_{n}*}|\right)\cdot\delta x^{\gamma}_{n}-\delta\log|g^{\gamma}_{b_{n}*}|\middle|x_{N}\right].\end{split}$
where $\delta x^{\gamma}_{n}=-g^{-1}_{b_{n}*}\delta g^{\gamma}_{b_{n}}(x_{n}).$
Proof.
By Section˜3.1, denote $\delta x^{\gamma}_{N-1}:=-g^{-1}_{b_{N-1}*}\delta g^{\gamma}_{b_{N-1}}(x_{N-1})$,
$\begin{split}\delta\log h^{\gamma}_{N}(x_{N})=\mathbb{E}\left[\delta\log h^{\gamma}_{N-1}+\nabla_{\delta x^{\gamma}_{N-1}}\log h_{N-1}-\delta\log|g^{\gamma}_{b_{N-1}*}|-\nabla_{\delta x^{\gamma}_{N-1}}\log|g_{b_{N-1}*}|\middle|x_{N}\right],\end{split}$
where all quantities are evaluated at $x_{N-1}$.
To handle the term involving $\nabla\log h_{N-1}$, note that it does not depend on $x_{N}$, so
$\begin{split}\mathbb{E}\left[\nabla_{\delta x^{\gamma}_{N-1}}\log h_{N-1}(x_{N-1})\middle|x_{N}\right]=\mathbb{E}\left[\mathbb{E}\left[\nabla\log h_{N-1}(x_{N-1})\middle|x_{N-1},x_{N}\right]\cdot\delta x^{\gamma}_{N-1}\middle|x_{N}\right].\end{split}$
use Definition˜1 (note that it requires conditioning on all steps $\geq N-1$),
$\begin{split}\mathbb{E}\left[\nabla\log h_{N-1}(x_{N-1})\middle|x_{N-1},x_{N}\right]=\mathbb{E}\left[\nu_{N-1}\middle|x_{N-1},x_{N}\right].\end{split}$
Since $\delta x^{\gamma}_{N-1}$ is measurable with respect to $(x_{N-1},x_{N})$,
$\begin{split}\mathbb{E}\left[\nabla_{\delta x^{\gamma}_{N-1}}\log h_{N-1}(x_{N-1})\middle|x_{N}\right]=\mathbb{E}\left[\mathbb{E}\left[\nu_{N-1}\cdot\delta x^{\gamma}_{N-1}\middle|x_{N-1},x_{N}\right]\middle|x_{N}\right]=\mathbb{E}\left[\nu_{N-1}\cdot\delta x^{\gamma}_{N-1}\middle|x_{N}\right].\end{split}$
By substitution,
$\begin{split}\delta\log h^{\gamma}_{N}(x_{N})=\mathbb{E}\left[\delta\log h^{\gamma}_{N-1}-\delta\log|g^{\gamma}_{b_{N-1}*}|+\left(\nu_{N-1}-\nabla\log|g_{b_{N-1}*}|\right)\cdot\delta x^{\gamma}_{N-1}\middle|x_{N}\right].\end{split}$
Repeat this on $\delta\log h^{\gamma}_{n}(x_{n})$ for each $n$ to prove the lemma. ∎
Note that we need the score of $h_{0}$, or $\nu_{0}$. Only for large $T$ and systems with fast decorrelation, we can use any $\nu_{0}$ and the impact would be negligible. See ˜Remark for more details.
3.3. Preparations for SDEs
This subsection introduces some notation for time-discretized SDEs and recalls some useful facts from [24] about scores for SDEs. We want to formally pass the discrete-time results to the continuous-time limit SDE,
(5) $\begin{split}dx^{\gamma}_{t}=F^{\gamma}_{t}(x^{\gamma}_{t})dt+\sigma^{\gamma}_{t}(x^{\gamma}_{t})dB_{t},\end{split}$
where $B$ denotes a standard Brownian motion. All SDEs in this paper are in the Ito sense, so the discretized version takes the form of Equation˜6. Denote the density of $x_{t}$ by $h_{t}$. Let $\mathcal{F}_{t}$ be the $\sigma$-algebra generated by $\{x_{\tau}\}_{\tau\leq t}$.
Our derivation is performed on the time span divided into small segments of length $\Delta t$}. Let $N$ be the total number of segments, so $N{\Delta t}=T$. Denote
${\Delta B}_{n}:=B_{n+1}-B_{n}.$
Denote $(\cdot)_{n}=(\cdot)_{n{\Delta t}}$. The time-discretized SDE is
(6) $x^{\gamma}_{n+1}-x^{\gamma}_{n}=F^{\gamma}(x^{\gamma}_{n}){\Delta t}+\sigma^{\gamma}(x^{\gamma}_{n}){\Delta B}_{n}.$
This section derives score formulas for time-discretized SDEs, then formally pass to the continuous-time limit. Comparing the notations, we have
(7) $\begin{split}f^{\gamma}_{n}(x):=x+F^{\gamma}_{n}(x){\Delta t},\quad b_{n}:={\Delta B}_{n},\quad\\ k(b):=\left(2\pi\sqrt{\Delta t}\right)^{-M/2}\exp\left(-\frac{\left\langle b,b\right\rangle}{2{\Delta t}}\right),\quad\nabla\log k\left({\Delta B}_{n}\right)=-\frac{{\Delta B}_{n}}{{\Delta t}},\\ g^{\gamma}_{b_{n}}(x_{n}):=x_{n}+F^{\gamma}_{n}(x_{n}){\Delta t}+\sigma^{\gamma}_{n}(x_{n}){\Delta B}_{n},\\ g^{\gamma}_{b_{n}*}(x_{n})=I+\nabla F^{\gamma}_{n}(x_{n}){\Delta t}+{\Delta B}_{n}\nabla\sigma_{n}^{\gamma T}(x_{n}).\end{split}$
Here $\nabla F$ and ${\Delta B}\nabla\sigma^{T}$ are matrices, whose $i$-th row $j$-th column are
$\begin{split}\left[\nabla F\right]_{ij}=\frac{\partial F^{i}}{\partial x^{j}}=:\partial_{j}F^{i},\quad\quad\left[{\Delta B}\nabla\sigma^{T}\right]_{ij}=\frac{\partial\sigma}{\partial x^{j}}{\Delta B}^{i}=:\partial_{j}\sigma{\Delta B}^{i}.\end{split}$
Ignoring terms that go to zero in the continuous-time limit, [24] derived explicit expressions of some terms for time-discretized SDEs,
(8) $\begin{split}\left|g^{\gamma}_{b_{n}*}\right|\approx 1+\sum_{i=1}^{M}\left(\partial_{i}F_{n}^{\gamma i}{\Delta t}+\partial_{i}\sigma_{n}^{\gamma}{\Delta B}^{i}\right)=1+\textnormal{div}F_{n}^{\gamma}{\Delta t}+\nabla\sigma_{n}^{\gamma}\cdot{\Delta B},\\ \textnormal{div}g_{b_{n}*}(x_{n}):=\nabla\log\left|g_{b*}\right|\approx\nabla\textnormal{div}F_{n}(x_{n}){\Delta t}+\nabla^{2}\sigma_{n}(x_{n}){\Delta B}-\nabla^{2}\sigma_{n}\nabla\sigma_{n}(x_{n}){\Delta t},\end{split}$
where $\nabla^{2}\sigma{\Delta B}=\sum_{i}\nabla\partial_{i}\sigma(x_{n}){\Delta B}_{n}^{i}$, $\nabla^{2}\sigma\nabla\sigma{\Delta t}=\sum_{i}\nabla\partial_{i}\sigma(x_{n})\partial_{i}\sigma(x_{n}){\Delta t}$. This paper also needs expressions of some other terms, which will be given in Section˜3.4. Then we recall the expression of the score function at $\gamma=0$.
{lemma}
[formal divergence-kernel fomula for score of SDEs [24]] For the SDE in $\mathbb{R}^{M}$,
$dx_{t}=F_{t}(x_{t})dt+\sigma_{t}(x_{t})dB_{t},$
For any backward adapted process $\{\alpha_{t}\}_{t=0}^{T}$, let $\{\nu_{t}\}_{t=0}^{T}$ be the forward covector process
$\begin{split}d\nu=\left((\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha)\nu-\nabla\textnormal{div}F+\nabla^{2}\sigma\nabla\sigma+\nabla\sigma\Delta\sigma\right)dt-(\nabla\sigma\nu^{T}+\nabla^{2}\sigma+\alpha\sigma^{-1})dB,\end{split}$
with initial condition $\nu_{0}=\nabla\log h_{0}(x_{0})$, where $h_{t}$ is the probability density of $x_{t}$, $\Delta\sigma$ is the Laplacian of the scalar-valued function $\sigma$. Then
$\begin{split}\nabla\log h_{T}(x_{T})=\mathbb{E}\left[\nu_{T}\middle|x_{T}\right].\end{split}$
Here, the SDE for $\nu$ is the limit of
(9) $\begin{split}\Delta\nu_{n}\approx\left((\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha_{n+1})\nu-\nabla\textnormal{div}F+\nabla^{2}\sigma\nabla\sigma+\nabla\sigma\Delta\sigma\right){\Delta t}\\ -(\nabla\sigma\nu^{T}+\nabla^{2}\sigma+\alpha_{n+1}\sigma^{-1}){\Delta B}\end{split}$
as ${\Delta t}\rightarrow 0$. Here $\Delta\nu_{n}:=\nu_{n+1}-\nu_{n}$; $\alpha_{n+1}$ is evaluated at step $n+1$; all other terms are evaluated at step $n$ and location $x_{n}$.
3.4. Continuous-time divergence-kernel formula for linear response
First, we give an expression of $\delta\log|g^{\gamma}_{b_{n}*}|$, then we derive the linear response formulas.
{lemma}
Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$,
$\begin{split}\delta\log|g^{\gamma}_{b_{n}*}|\approx\textnormal{div}\delta F_{n}^{\gamma}{\Delta t}+\nabla\delta\sigma_{n}^{\gamma}\cdot{\Delta B}_{n}-\nabla\sigma_{n}^{T}\nabla\delta\sigma_{n}^{\gamma}{\Delta t}.\end{split}$
Proof.
Substitute the expression in Equation˜8, we get
$\begin{split}\delta\log|g^{\gamma}_{b_{n}*}|=\frac{\delta\left|g^{\gamma}_{b*}\right|}{\left|g_{b*}\right|}=\frac{\delta\textnormal{div}F^{\gamma}{\Delta t}+\delta\nabla\sigma^{\gamma}\cdot{\Delta B}}{1+\textnormal{div}F{\Delta t}+\nabla\sigma\cdot{\Delta B}}.\end{split}$
Recall that $1/(1+\varepsilon)=1-\varepsilon+O(\varepsilon^{2})$, ignoring high-order terms, we get
$\begin{split}\delta\log|g^{\gamma}_{b_{n}*}|\approx\left(\textnormal{div}\delta F^{\gamma}{\Delta t}+\nabla\delta\sigma^{\gamma}\cdot{\Delta B}\right)\left({1-\textnormal{div}F{\Delta t}-\nabla\sigma\cdot{\Delta B}}\right)\\ \approx\textnormal{div}\delta F^{\gamma}{\Delta t}+\nabla\delta\sigma^{\gamma}\cdot{\Delta B}-(\nabla\delta\sigma^{\gamma}\cdot{\Delta B})(\nabla\sigma\cdot{\Delta B})\end{split}$
For the last term, ignore cross terms (${\Delta B}^{i}{\Delta B}^{j}$ with $i\neq j$) and apply ${\Delta B}^{i}{\Delta B}^{i}\approx{\Delta t}$,
$\begin{split}(\nabla\delta\sigma^{\gamma}\cdot{\Delta B})(\nabla\sigma\cdot{\Delta B})=\sum_{i}\sum_{j}\partial_{i}\delta\sigma^{\gamma}{\Delta B}^{i}\partial_{j}\sigma{\Delta B}^{j}\approx\sum_{i}\partial_{i}\delta\sigma^{\gamma}\partial_{i}\sigma{\Delta t}=\nabla\sigma^{T}\nabla\delta\sigma^{\gamma}{\Delta t}.\end{split}$
Then substitute to prove the corollary. ∎
{lemma}
[expression of $\delta x^{\gamma}_{n}$] Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$,
$\begin{split}\delta x^{\gamma}_{n}=-g^{-1}_{b_{n}*}\delta g^{\gamma}_{b_{n}}(x_{n})\approx-\delta F^{\gamma}_{n}{\Delta t}-\delta\sigma^{\gamma}_{n}{\Delta B}_{n}+\delta\sigma^{\gamma}_{n}\nabla\sigma_{n}{\Delta t}.\end{split}$
Remark.
This is one of the examples where some terms in the SDE setting are more explicit than discrete-time systems. For a discrete-time system, this term would involve solving a $M$-dimensional linear equation system.
Proof.
By Equation˜7,
$\begin{split}g^{\gamma}_{b_{n}}(x_{n}):=x_{n}+F^{\gamma}_{n}(x_{n}){\Delta t}+\sigma^{\gamma}_{n}(x_{n}){\Delta B}_{n},\\ g_{b_{n}*}(x_{n})=I+\nabla F_{n}(x_{n}){\Delta t}+{\Delta B}_{n}\nabla\sigma_{n}^{T}(x_{n}).\end{split}$
Hence,
$\begin{split}\delta g^{\gamma}_{b_{n}}(x_{n})=\delta F^{\gamma}_{n}(x_{n}){\Delta t}+\delta\sigma^{\gamma}_{n}(x_{n}){\Delta B}_{n}\sim O({\Delta t}^{0.5}).\end{split}$
For a small matrix $\varepsilon\sim O({\Delta t}^{0.5})$, $(I+\varepsilon)^{-1}=I-\varepsilon+O({\Delta t})$, so
$\begin{split}g_{b_{n}*}^{-1}\approx I-\nabla F_{n}{\Delta t}-{\Delta B}_{n}\nabla\sigma_{n}^{T}.\end{split}$
Hence, ignoring $O({\Delta t}^{1.5})$ terms which vanish as ${\Delta t}\rightarrow 0$,
$\begin{split}\delta x^{\gamma}_{n}=-g_{b_{n}*}^{-1}\delta g^{\gamma}_{b_{n}}(x_{n})\approx-(I-\nabla F_{n}{\Delta t}-{\Delta B}_{n}\nabla\sigma^{T}_{n})(\delta F^{\gamma}_{n}{\Delta t}+\delta\sigma^{\gamma}_{n}{\Delta B}_{n})\\ \approx-\delta F^{\gamma}_{n}{\Delta t}-\delta\sigma^{\gamma}_{n}{\Delta B}_{n}+{\Delta B}_{n}\nabla\sigma^{T}_{n}\delta\sigma^{\gamma}_{n}{\Delta B}_{n}\end{split}$
Note that $\nabla\sigma^{T}_{n}{\Delta B}_{n}$ is a number, and ignoring the cross terms, we get
$\begin{split}{\Delta B}_{n}(\nabla\sigma^{T}_{n}{\Delta B}_{n})={\Delta B}_{n}({\Delta B}_{n}^{T}\nabla\sigma_{n})\approx\nabla\sigma_{n}{\Delta t}.\end{split}$
Then substitute to prove the lemma. ∎
{lemma}
Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$, let $\Delta\sigma$ denote the Laplacian of $\sigma$, then
$\begin{split}\nabla\log\left|g_{b_{n}*}\right|\cdot\delta x^{\gamma}_{n}\approx-\delta\sigma^{\gamma}_{n}\Delta\sigma_{n}{\Delta t}.\end{split}$
Proof.
By Equation˜8 and Section˜3.4,
$\begin{split}\nabla\log\left|g_{b_{n}*}\right|\approx\nabla\textnormal{div}F_{n}(x_{n}){\Delta t}+\nabla^{2}\sigma_{n}(x_{n}){\Delta B}_{n}-\nabla^{2}\sigma_{n}\nabla\sigma_{n}(x_{n}){\Delta t},\\ \delta x^{\gamma}_{n}=-g^{-1}_{b_{n}*}\delta g^{\gamma}_{b_{n}}(x_{n})\approx-\delta F^{\gamma}_{n}{\Delta t}-\delta\sigma^{\gamma}_{n}{\Delta B}_{n}+\delta\sigma^{\gamma}_{n}\nabla\sigma^{T}_{n}{\Delta t}.\end{split}$
In the product, the only term smaller than or equal to $O({\Delta t})$ is
$\begin{split}\nabla\log\left|g_{b_{n}*}\right|\cdot\delta x^{\gamma}_{n}\approx\nabla^{2}\sigma_{n}(x_{n}){\Delta B}_{n}\cdot-\delta\sigma^{\gamma}_{n}{\Delta B}_{n}\end{split}$
Ignore the cross term, we get
$\begin{split}\nabla^{2}\sigma{\Delta B}\cdot{\Delta B}=\sum_{i,j}\partial_{i}\partial_{j}\sigma{\Delta B}^{j}{\Delta B}^{i}\approx\sum_{i}\partial_{i}^{2}\sigma{\Delta t}=:\Delta\sigma{\Delta t},\end{split}$
Then substitute to prove the lemma. ∎
With these, we can formally pass Definition˜1 to the continuous-time limit.
See 1.1
Proof.
Substitute Sections˜3.4, 3.4 and 3.4 into Definition˜1, we get
(10) $\begin{split}\delta\log h^{\gamma}_{N}(x_{N})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}+\sum_{n=0}^{N-1}\nu_{n}\cdot\delta x^{\gamma}_{n}-\nabla\log|g_{b_{n}*}|\cdot\delta x^{\gamma}_{n}-\delta\log|g^{\gamma}_{b_{n}*}|\middle|x_{N}\right]\\ \approx\mathbb{E}\left[\delta\log h^{\gamma}_{0}+\sum_{n=0}^{N-1}\nu_{n}\cdot(-\delta F^{\gamma}_{n}{\Delta t}-\delta\sigma^{\gamma}_{n}{\Delta B}_{n}+\delta\sigma^{\gamma}_{n}\nabla\sigma_{n}{\Delta t})\right.\\ \left.+\delta\sigma^{\gamma}_{n}\Delta\sigma_{n}{\Delta t}-\textnormal{div}\delta F_{n}^{\gamma}{\Delta t}-\nabla\delta\sigma_{n}^{\gamma}\cdot{\Delta B}_{n}+\nabla\sigma_{n}^{T}\nabla\delta\sigma_{n}^{\gamma}{\Delta t}\middle|x_{N}\right].\end{split}$
Then formally pass to the limit ${\Delta t}\rightarrow 0$. ∎
We also formally pass to the infinite-time limit of stationary distributions $h^{\gamma}$. We need to further assume that $F_{t}$ and $\sigma_{t}$ are stationary, and that there is $T^{\prime}$ such that $\alpha_{t}$ only depends on $x_{t}\sim x_{T^{\prime}}$, so $(\alpha,\nu,x)$ are jointly stationary. {corollary} [formal ergodic divergence-kernel linear response] For any $C^{2}$ observable function $\Phi:\mathbb{R}^{M}\rightarrow R$, let $h^{\gamma}$ be the stationary measure under parameter $\gamma$, then the linear response of the averaged observable can be expressed almost surely on one path:
[…]
The discrete-time expression is
$\begin{split}\delta\int\Phi(x)h^{\gamma}(x)dx\overset{a.s.}{=}\lim_{W\rightarrow\infty}\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=0}^{N}\left(\sum_{m=1}^{W}(\Phi(x_{n+m})-\Phi_{avg})\right)\nu_{n}\cdot\delta\sigma^{\gamma}_{n}\nabla\sigma_{n}{\Delta t}\\ -\nu_{n}\cdot(\delta F^{\gamma}_{n}{\Delta t}+\delta\sigma^{\gamma}_{n}{\Delta B}_{n})+\delta\sigma^{\gamma}_{n}\Delta\sigma_{n}{\Delta t}-\textnormal{div}\delta F_{n}^{\gamma}{\Delta t}-\nabla\delta\sigma_{n}^{\gamma}\cdot{\Delta B}_{n}+\nabla\sigma_{n}^{T}\nabla\delta\sigma_{n}^{\gamma}{\Delta t}.\end{split}$
3.5. Discussions
The discussions in [24] are still valid. Roughly speaking, we want $\alpha$ larger than the largest contraction rate. Also, the divergence-kernel method does not work well in highly contracting systems with little to no noise, the case where the path-perturbation method is good. We may need to use all three methods (path, kernel, and divergence) together to obtain an ultimate solution, as proposed in [25].
Compared to the path-kernel method in [23, 22], the divergence-kernel method has the following advantages:
  1. (1)
    It involves only forward processes, so the memory cost is much smaller.
  2. (2)
    Its cost is almost independent of both the number of observables and the parameters. In contrast, the cost of the path-kernel method is independent of either the number of observables or the parameters.
It also has some disadvantages:
  1. (1)
    Its computational complexity per step is higher than that for the path-kernel method, although still on the same order as that for the path-kernel method. This is because it involves the second-order derivative.
  2. (2)
    For dynamical systems generated from discretized PDEs, typically $|\lambda_{min}|>|\lambda_{max}|$, where $\lambda_{max}$ and $\lambda_{min}$ are the largest and the smallest (most negative) Lyapunov exponents; hence, the path-kernel method requires a smaller $\alpha$, so the estimator has smaller size and the average converges faster.
4. Numerical examples
4.1. 1-dimensional process with multiplicative noise
We use Section˜1.1 to compute the linear response of the one-dimensional process with multiplicative noise. The SDE is
$\begin{split}dx=\beta_{t}(\gamma-x)dt+\sqrt{\beta_{t}}(0.5+\exp\left((\gamma-x)^{2}\right)dB,\quad\textnormal{where}\quad\beta_{t}:=1+3t.\end{split}$
The initial distribution is $x_{0}\sim N(0,1)$. Set $T=0.1$, ${\Delta t}=0.01$, $\alpha=10$. Note that the problem is symmetric for $\gamma$.
We are only interested in terminal conditions $x_{T}$ in the interval [2,2][-2,2], which is further partitioned into $10$ subintervals. We simulate many sample paths, then count the number of paths whose $x_{T}$ fall into a specific subinterval, to obtain the empirical distribution of $x_{T}$ on subintervals. Then we compute the linear response $\delta\log h_{T}$ by averaging the divergence-kernel formula over these paths, and see if our results reflect the trend of $\log h_{T}$ from the empirical distribution. We also use the average of $\nu_{T}$ to compute the scores $\nabla\log h_{T}$. Each subinterval contains at least 180 samples. The $\delta\log h_{T}$ and $\nabla\log h_{T}$ computed for different $\gamma$ and $x$’s are shown in Figure˜1. Both correctly reflect the trend of $\log h_{T}$.
Figure 1. Divergence-kernel method for the linear responses and scores of the 1-dimensional SDE with multiplicative noise, $T=0.3$. The dots are $\log h_{T}$. Each short line is a linear response or score computed by the divergence-kernel algorithm, averaged on paths whose terminal conditions $x_{T}$ fall into the same subinterval. Left: linear responses $\delta\log h_{T}$. Right: scores $\nabla\log h_{T}$ .
4.2. 40-dimensional Lorenz 96 system with multiplicative noise
We use ˜Remark to compute the linear response of the stationary density of the Lorenz 96 model [14]. The dimension of the system is $M=40$. The SDE is
$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+8+\gamma-0.01(x^{i})^{2}\right)dt+0.5+\exp\left((x-\gamma)^{2}\right)dB^{i},$
$\textnormal{where}\quad i=1,\ldots,M;\quad x_{0}\sim{\mathcal{N}}(0,I).$
Here $i$ labels different directions in $\mathbb{R}^{M}$, and it is assumed that $x^{-1}=x^{M-1},x^{0}=x^{M}$ and $x^{M+1}=x^{1}$. We added noise and the $0.01(x^{i})^{2}$ term, which prevents the noise from carrying us to infinitely far away. We choose the observable function to be
$\Phi(x):=\frac{1}{M}|x|^{2}.$
We use the Euler integration scheme with ${\Delta t}=0.002$. A typical orbit is in Figure˜2. Then we use the ergodic linear response formula ˜Remark to compute the linear response of the observable averaged on the stationary measure. The averaged observable $\Phi_{avg}$ and the linear response $\delta\int\Phi h^{\gamma}$ are averaged over 7 orbits of length $T=400$. The decorrelation time is $W=1.5$. The results shown in Figure˜2; the linear response computed using the divergence-kernel method is correct.
Figure 2. Left: plot of $x^{0}_{t},x^{1}_{t}$ from a typical orbit of time length $1.5$. Right: linear response computed by divergence-kernel method.
5. A simple forward-only diffusion model
We consider the problem in which we are given $N_{data}$ many samples $y_{k}\in\mathbb{R}^{M}$, whose true distribution is unknown. Then we use gradient descent to find an SDE such that it maps the Gaussian distribution to match the data.
First, we describe how to generate samples. Note that the information in this paragraph will not be available when we infer the SDE. The samples are generated by the SDE
$\begin{split}dx_{t}=F^{\gamma_{true}}_{t}(x_{t})dt+\sigma^{\gamma_{true}}_{t}(x_{t})dB_{t},\end{split}$
where the true parameters $\gamma_{true}:=[\gamma_{true}^{0},\ldots,\gamma_{true}^{N_{\gamma}}]$. The samples are $y:=x_{T}$ obtained at the end of independent paths. Denote the empirical measure, which we try to match, as
$\begin{split}P(\cdot):=\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\delta_{y_{k}}(\cdot).\end{split}$
Our framework can work on more general data, but here we restrict ourselves to the simple case, so that after optimization, we can check our result with the true parameters.
We want to fit several unknown parameters $\gamma=[\gamma^{0},\ldots,\gamma^{N_{\gamma}}]$ in an SDE such that its marginal distribution matches the data. The prototype of the SDE is
$\begin{split}dx_{t}=F^{\gamma}_{t}(x_{t})dt+\sigma^{\gamma}_{t}(x_{t})dB_{t}.\end{split}$
Denote the density at $T$ by $h^{\gamma}$ and measure by $H^{\gamma}$ – this is the distribution that we want to optimize. For each $\gamma$ during optimization, we simulate $L$-many paths by this SDE, and we label each path by $l$, in particular, the terminal location is denoted by $x_{T,l}$. In future research, we can seek to use more flexible prototypes, such as neural networks. However, note that the gradient used in the optimization is not computed by backpropagation; rather, it only involves a covector process that runs forward in time and costs much less memory. We also remark that we can choose to optimize the stationary distribution using ˜Remark, which may be more efficient for some cases.
The difference between the generated distribution and the data is measured by the KL-divergence
$\begin{split}D_{\text{KL}}(P\parallel H^{\gamma}):=\int_{y\in{\mathbb{R}^{M}}}\log{\frac{P(dy)}{H^{\gamma}(dy)}}\,P(dy).\end{split}$
Its parameter-derivative is
$\begin{split}\delta D_{\text{KL}}(P\parallel H^{\gamma})=-\int_{y\in{\mathbb{R}^{M}}}\frac{\delta H^{\gamma}(dy)}{H^{\gamma}(dy)}\,P(dy)=-\int_{y\in{\mathbb{R}^{M}}}\frac{\delta h^{\gamma}}{h^{\gamma}}(y)\,P(dy)\\ =-\int_{y\in{\mathbb{R}^{M}}}\delta\log h^{\gamma}(y)\,P(dy)=-\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\delta\log h^{\gamma}(y_{k}).\end{split}$
With the new formula in Section˜1.1, we have an expression of $\delta\log h^{\gamma}$ at $x_{T,l}$’s, which enables the gradient optimization on this KL divergence.
Then we need to decide how to compute $\delta\log h^{\gamma}$ at the prescribed locations $y_{k}$’s. It is difficult to obtain paths ending exactly at $y_{k}$, so we need an approximation. We approximate $y_{k}$ by its $N_{neighbor}$-nearest neighbors in $\{x_{T,l}\}_{l=1}^{L}$, and therefore $\delta\log h^{\gamma}(y_{k})$ is approximated by the average of the divergence-kernel expression in Section˜1.1 over those $N_{neighbor}$ paths. Hence,
$\begin{split}\delta D_{\text{KL}}(P\parallel H^{\gamma})\approx\hat{\delta}D_{\text{KL}}(P\parallel H^{\gamma}):=-\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\frac{1}{N_{neighbor}}\sum_{x_{T,l}\in N(y_{k})}\mathcal{T}(x_{T,l}),\end{split}$
where $N(y_{k})$ is the neighbors of $y_{k}$, and $\mathcal{T}(x_{T,l})$ is the expression in Section˜1.1 on the path ending at $x_{T,l}$. After approximately computing the parameter-derivative, we update the parameter by gradient descent,
$\begin{split}\gamma_{n+1}=\gamma_{n}-\eta\hat{\delta}D_{\text{KL}}(P\parallel H^{\gamma}).\end{split}$
For simplicity, we keep $\eta$ constant, but we can also change it after each update of $\gamma$.
5.1. A 1-d problem with 2 parameters
First, we try a 1-dimensional problem. The SDE is
$\begin{split}dx^{i}=\left(\gamma^{0}-x\right)dt+\left(0.5+\exp\left(-(x-\gamma^{1})^{2}\right)\right)dB,\quad\textnormal{where}\quad x_{0}\sim{\mathcal{N}}(0,1).\end{split}$
The true parameter used to generate the data is $\gamma_{true}=[0,0]$. The initial parameter in the optimization is $\gamma_{0}=[5,1]$. The number of data points is $N_{data}=200$. The number of samples generated for each update of $\gamma$ is $L=200$. The stepsize (or learning rate) is $\eta=1$. The number of neighbors used to approximate each $y_{k}$ is $N_{neighbor}=5$.
We start from $\gamma_{0}$ and use gradient descent to minimize the KL-divergence. The histograms of $y_{k}$’s and $x_{T,l}$’s for $\gamma_{0}$ and $\gamma_{10}$, the parameter after 10 updates, are shown in Figure˜3. As we can see, the two distributions get much closer during the optimization. More specifically, the error $\gamma-\gamma_{true}$ is reduced from $[5,1]$ to [0.05,0.09][-0.05,0.09]. The computational time for all 10 updates on a 4-core CPU is about 2 minutes, but note that our implementation is far from optimal.
Figure 3. Histograms of $y_{k}$’s and $x_{T,l}$’s. Left: plot at $\gamma_{0}$. Right: at $\gamma_{10}$.
5.2. A 5-d problem with 6 parameters
We try a 5-dimensional Lorenz 96 system. The SDE is
$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+\gamma^{i}-0.01(x^{i})^{2}\right)dt+\left(0.5+\exp\left((x-\gamma^{M}\mathbbm{1})^{2}\right)\right)dB^{i},$
$\textnormal{where}\quad i=0,\ldots,M-1;\quad M=5;\quad\mathbbm{1}=[1,1,1,1,1];\quad x_{0}\sim{\mathcal{N}}(0,I).$
The true parameter used to generate the data is $\gamma_{true}=[5,6,7,8,9,2]$. The initial parameter in the optimization is $\gamma_{0}=[0,0,0,0,0,0]$. The number of data points is $N_{data}=200$. The number of samples generated for each update of $\gamma$ is $L=200$. The stepsize (or learning rate) is $\eta=1$. The number of neighbors used to approximate each $y_{k}$ is $N_{neighbor}=5$.
We start from $\gamma_{0}$ and use gradient descent to minimize the KL-divergence. The history of the $l^{2}$-distance $|\gamma_{n}-\gamma_{true}|$ for $n=0$ to $50$ is shown in Figure˜4. As we can see, $\gamma_{n}$ gets closer to $\gamma_{true}$ during optimization, with $\gamma_{48}$ achieving the minimum distance. More specifically,
$\begin{split}\gamma_{0}-\gamma_{true}=[-5,-6,-7,-8,-9,-2];\\ \gamma_{48}-\gamma_{true}=[-0.33,0.40,-0.26,-0.20,-0.24,-0.36].\end{split}$
The computational time for all 50 updates on a 4-core CPU is about 2 minutes, but note that our implementation is far from optimal.
Figure 4. History of $|\gamma_{n}-\gamma_{true}|$ for the 5-d diffusion model.
Declarations
Data availability statement
The code used in this paper is posted at https://github.com/niangxiu/DKlinR. There are no other associated data.
Funding
The author has no active funding.
Conflicts of interests/Competing interests
The author has no relevant conflicting or competing interests.
Appendix A Kernel-differentiation method
We briefly go through the kernel-differentiation method, and explain its problem when applied to linear responses of high-dimensional systems with respect to diffusion coefficients.
A.1. One-step results
Let $x^{\gamma}_{0}\sim h^{\gamma}_{0}$, $b^{\gamma}_{0}\sim k$, the one-step dynamics is
$\begin{split}x^{\gamma}_{1}=f^{\gamma}_{0}(x^{\gamma}_{0})+\sigma^{\gamma}_{0}(x^{\gamma}_{0})b_{0}.\end{split}$
The most natural expression of $h_{1}$ is:
$h^{\gamma}_{1}(x_{1})=\int h^{\gamma}_{0}(x_{0})(\sigma^{\gamma}_{0})^{-M}(x_{0})k\left(\frac{x_{1}-f^{\gamma}_{0}(x_{0})}{\sigma^{\gamma}_{0}(x_{0})}\right)dx_{0}.$
Take derivative with respect to $\gamma$, then the probability kernel $k$ will be differentiated, and we get the kernel-differentiation formula for one-step.
{lemma}
[Kernel formula for one-step linear response]
$\delta\log h^{\gamma}_{1}(x_{1})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}(x_{0})-M\delta\log\sigma^{\gamma}_{0}(x_{0})-\frac{\nabla\log k(b_{0})}{\sigma_{0}(x_{0})}\left(\delta\sigma^{\gamma}_{0}(x_{0})b_{0}+\delta f^{\gamma}(x_{0})\right)\middle|x_{1}\right].$
Proof.
Take derivative of Equation˜2 with respect to $\gamma$,
$\begin{split}\delta h^{\gamma}_{1}(x_{1})=\int\left(\frac{\delta h^{\gamma}_{0}}{h_{0}}(x_{0})-M\frac{\delta\sigma^{\gamma}_{0}}{\sigma_{0}}(x_{0})-\frac{\nabla k}{k}(b_{0})\left(\frac{\delta\sigma^{\gamma}_{0}}{\sigma_{0}}(x_{0})b_{0}+\frac{\delta f^{\gamma}_{0}}{\sigma_{0}}(x_{0})\right)\right)h_{0}(x_{0})\frac{k\left(b_{0}\right)}{\sigma^{M}(x_{0})}dx_{0},\end{split}$
where $b_{0}=(x_{1}-f_{0}(x_{0}))/\sigma_{0}(x_{0})$. Divide both sides by $h_{1}(x_{1})$ and apply the definition of the conditional expectation to prove the lemma. ∎
The second term in Section˜A.1 can be very large for high-dimensional noise since it involves $M$. For now, we still use noise in all directions, so we try to avoid using the kernel formula for parameter derivative. The kernel formula for the spatial derivative does not have terms involving $M$ [24], so we can still use that. We will try to use low-dimensional noise in later papers, which will incur additional difficulties.
A.2. Discrete-time many-steps
First by recursively apply Section˜A.1, we can prove the following kernel-differentiation formula for linear response. Here, if we set $\delta\sigma=0$, this becomes the typical likelihood ratio formula.
{lemma}
[N-step kernel-differentiation formula for linear responses]
$\delta\log h^{\gamma}_{N}(x_{N})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}(x_{0})-\sum_{n=0}^{N-1}\frac{1}{\sigma_{n}(x_{n})}\left(M\delta\sigma^{\gamma}_{n}(x_{n})+\nabla\log k(b_{n})\left(\delta\sigma^{\gamma}_{n}(x_{n})b_{n}+\delta f^{\gamma}(x_{n})\right)\right)\middle|x_{N}\right].$
A.3. Formal continuous-time limit
For the continuous time limit, we first consider the time-discretized SDEs. Note that if $\delta\sigma\neq 0$, then the two terms involve $\delta\sigma$ in the kernel formula in Section˜A.1 becomes $O(1)$ in each step, so the integration in time does not converge – this is why likelihood ratio method always uses additive noise independent of the parameter $\gamma$. For the degenerate case where $\delta\sigma\equiv 0$, Section˜A.2 converges to the classical likelihood ratio method in continuous time.
{lemma}
[likelihood ratio formula for linear response of SDEs] If $\delta\sigma\equiv 0$,
$\begin{split}\nabla\log h_{T}(x_{T})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}(x_{0})+\int_{t=0}^{T}\frac{1}{\sigma_{t}(x_{t})}\delta F^{\gamma}(x_{n})\cdot dB\middle|x_{N}\right].\end{split}$

### v2

@@
-Divergence-Kernel method for linear responses and diffusion models
+Divergence-kernel method for linear responses of densities and generative models
 Abstract.
-We derive the divergence-kernel formula for the linear response (parameter-derivative of marginal or stationary distributions) of random dynamical systems, and formally pass to the continuous-time limit.
+We derive the divergence-kernel formula for the linear response of random dynamical systems.
+Specifically, the pathwise expression is for the parameter-derivative of the marginal or stationary density, not an averaged observable.
 Our formula works for multiplicative and parameterized noise over any period of time; it does not require hyperbolicity.
-Then we derive a pathwise Monte-Carlo algorithm for linear responses.
-With this, we propose a forward-only diffusion generative model and test on simple problems.
+Then we derive a Monte-Carlo algorithm for linear responses.
+We develop a new framework of generative models, DK-SDE, where the model is a parameterized SDE, that (1) directly uses the KL divergence between the empirical data distribution and the marginal density of the SDE as the training objective, and (2) accommodates parametrizations in both drift and diffusion over a long time span, allowing prior structural knowledge to be incorporated explicitly.
+The optimization is done by gradient-descent enabled by the divergence-kernel method, which involves only forward processes and therefore substantially reduces memory cost.
+We demonstrate the new model on a 20-dimensional Lorenz system.
 AMS subject classification numbers. 60H07, 60J60, 65D25, 65C30, 65C05.
-Keywords. Linear response, Divergence method, Likelihood ratio, SDE, Diffusion model.
+Keywords. Linear response, Divergence method, Likelihood ratio, SDE, Generative model.
 1. Introduction
 1.1. Main results
-This paper rigorously derives the divergence-kernel formula for the linear response of discrete-time random dynamical systems.
-Then we formally pass to the continuous-time limit.
-Our result gives a fundamental tool for the optimization of diffusion processes; in particular, from now on, we can use the gradient-based method to optimize marginal or stationary densities with respect to the diffusion coefficients.
-In particular, we propose a simple framework for diffusion generative models.
-Our main theoretical result is
-{theorem}
-[formal divergence-kernel formula for linear responses of SDEs] Consider the SDE in $\mathbb{R}^{M}$ parameterized by $\gamma$,
+This paper rigorously derives the divergence-kernel formula for the linear response in discrete-time random dynamical systems.
+It is a pathwise expression of the derivative of marginal or stationary density with respect to parameters in the drift or diffusion coefficients.
+Then we formally pass to the continuous-time limit in Section˜1.1.
+[…]
+Consider the SDE in $\mathbb{R}^{M}$ parameterized by $\gamma$,
 $dx^{\gamma}_{t}=F^{\gamma}_{t}(x^{\gamma}_{t})dt+\sigma^{\gamma}_{t}(x^{\gamma}_{t})dB_{t},$
 For any backward adapted process $\{\alpha_{t}\}_{t=0}^{T}$, let $\{\nu_{t}\}_{t=0}^{T}$ be the forward covector process at $\gamma=0$ such that
-$\begin{split}d\nu=\left((\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha)\nu-\nabla\textnormal{div}F+\nabla^{2}\sigma\nabla\sigma+\nabla\sigma\Delta\sigma\right)dt-(\nabla\sigma\nu^{T}+\nabla^{2}\sigma+\alpha\sigma^{-1})dB.\end{split}$
-with initial condition $\nu_{0}=\nabla\log h_{0}(x_{0})$, where $h_{t}$ is the probability density of $x_{t}$. Then $\nabla\log h_{t}(x_{t})=\mathbb{E}\left[\nu_{t}\middle|x_{t}\right]$, and
+$\begin{split}d\nu=\left((\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha)\nu-\nabla\textnormal{div}F+\nabla^{2}\sigma\nabla\sigma+\nabla\sigma\Delta\sigma\right)dt-(\nabla\sigma\nu^{T}+\nabla^{2}\sigma+\alpha\sigma^{-1})dB\end{split}$
+with initial condition $\nu_{0}=\nabla\log h_{0}(x_{0})$, where $h_{t}$ is the probability density of $x_{t}$.
+Then $\nabla\log h_{t}(x_{t})=\mathbb{E}\left[\nu_{t}\middle|x_{t}\right]$.
+Moreover, let $\delta(\cdot):=\partial(\cdot)/\partial\gamma$, then
 $\begin{split}\delta\log h^{\gamma}_{T}(x_{T})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}+\int_{t=0}^{T}\delta\sigma^{\gamma}\nabla\sigma^{T}\nu-\delta F^{\gamma T}\nu-\textnormal{div}\delta F^{\gamma}+\delta\sigma^{\gamma}\Delta\sigma+\nabla\sigma^{T}\nabla\delta\sigma^{\gamma}dt\right.\\ \left.-\int_{t=0}^{T}(\delta\sigma^{\gamma}\nu^{T}+\nabla\delta\sigma^{\gamma T})dB\middle|x_{T}\right].\end{split}$
 Here $\Delta\sigma$ is the Laplacian of $\sigma$, $B$ is the Brownian motion, and the SDE is Ito.
+In this paper $\sigma$ is assumed to be a scalar function uniformly bounded away from 0, and extension to matrices is conceptually straightforward.
 The time integrations are the limits of Equations˜9 and 10.
-More generally, $\nu$ is an estimator for the score of the SDE at $\gamma=0$, which can be replaced by other divergence-kernel expressions from [24].
-The main advantages of this result are:
+For large $T$, we can just pick $\alpha$ larger than the magnitude of the strongest contraction rate, which stabilizes $\nu$ from growing exponentially fast.
+In fact, for memory-light implementations we restrict to $\alpha_{t}=\alpha(x_{t})$ or constant $\alpha$; the backward-adapted freedom is mainly a theoretical gauge.
+The conditional expectation is taken over all paths ending at the same point; it turns pathwise quantities into terminal density derivative.
+The main advantages of this formula are:
   1. (1)
-    It computes derivative of the marginal density, not an averaged observable;
+    Compared with the path-kernel method in [30, 29], it computes derivative of the marginal density, not an averaged observable.
+Moreover, it involves forward processes so the memory cost is lower than the adjoint path-kernel method, yet it still computes the full gradient with cost independent of the number of parameters.
   2. (2)
-    The expression in pathwise, so we can compute by Monte-Carlo type algorithms in high-dimensions;
+    Compared with the finite-element method in [12], its main expression is pathwise, so we can compute it via Monte Carlo algorithms in high dimensions.
   3. (3)
-    The diffusion coefficient $\sigma$ can depend on the parameter $\gamma$ and the location $x$;
+    Compared with the likelihood ratio method in [15], the diffusion coefficient $\sigma$ can depend on the parameter $\gamma$ and the location $x$.
   4. (4)
-    The size of the estimator does not grow exponentially to $T$, so we can easily extend our result to stationary distributions (see ˜Remark);
+    Compared with the fast response method in [33], it does not require hyperbolicity (in the sense of dynamical systems).
   5. (5)
-    It does not require hyperbolicity (in the sense of dynamical systems).
-As a less obvious application, we use this result to construct a new diffusion generative model, where we directly optimize the marginal distribution to reduce its KL-divergence with respect to the data.
-Importantly, we can optimize the diffusion coefficients of multiplicative noise.
+    The size of the expression does not grow exponentially to $T$, so we can easily extend our result to stationary distributions (see Section˜3.4).
+  6. (6)
+    Compared with the author’s divergence-kernel method for scores in [31], this paper considers the parameter-gradients, which enables a new framework of parametric SDE generative models.
+This formula gives an efficient gradient tool for diffusion processes with not too small noise or not too strong contraction.
+In Section˜4.2, the corresponding Monte-Carlo algorithm computes the linear response of the marginal density of a Lorenz 96 system, which is difficult for other methods.
+With the new divergence-kernel tool, we develop a new framework for generative models.
+The logic is rather simple.
+Our model is an SDE, whose drift and diffusion depends on the state and (many) undetermined parameters.
+The loss function is the KL-divergence of the empirical distribution of the data with respect to the marginal distribution of the SDE.
+With the new gradient tool, we can just optimize parameters in the SDE model via stochastic gradient descent.
+This new framework combines several advantages of different current generative models, and also offers some new advantages:
+  1. (1)
+    Directly uses the KL divergence between the empirical data distribution and the marginal density of the SDE as the training objective.
+  2. (2)
+    Accommodates parametrizations in both drift and diffusion, so prior knowledge of the SDE model can be easily incorporated into the model and reduce the complexity (number of parameters) of the model.
+  3. (3)
+    Allows deep models (large $T$) without imposing normalizing requirements.
+  4. (4)
+    It only saves the current state. In particular, we do not have a backward procedure, so we do not save the entire path, substantially reducing the memory cost.
+  5. (5)
+    If we use the stationary measure of the SDE model in the framework, then the ergodic formula in Section˜3.4 further improves efficiency.
+Finally, we might achieve better overall efficiency in general situations; but this requires many more experiments to verify.
 1.2. Literature review
+1.2.1. Linear response methods
 The marginal or stationary distribution of a random dynamical system is of central interest in applied sciences.
@@
 It also includes the backpropagation method, which is the basic algorithm for machine learning.
-However, it is expensive for chaotic or unstable system; the work-around is to artificially reduce the size of the path-perturbation, such as shadowing or clipping methods [26, 29, 30, 8], but they all introduce systematic errors.
+However, it is expensive for chaotic or unstable systems; the work-around is to artificially reduce the size of the path-perturbation, such as shadowing or clipping methods [34, 37, 39, 11], but they all introduce systematic errors.
 Moreover, it computes the derivative of an averaged observable, but does not give the derivative of the probability density.
@@
 However, it is expensive for high-dimensional systems with contracting directions.
-The kernel-differentiation method works only for random systems.
-It is also known as the Cameron-Martin-Girsanov theorem [4, 17] or the likelihood ratio method [33, 31, 11].
+For SDEs, we can take the derivative of the Fokker-Planck equation, which involves the divergence of the drift and a second-order differential operator [18], which is hard to sample on paths.
+The kernel-differentiation method works only for random systems; as the name indicates, the derivative now hits the probability kernel.
+It is also known as the Cameron-Martin-Girsanov theorem [6, 25], the likelihood ratio method [42, 40, 15], or the Monte-Carlo gradient method [5].
 However, it cannot handle multiplicative noise or perturbation on the diffusion coefficients.
@@
 Our path-kernel method in [30] gives the pathwise linear response formula of the diffusion coefficients.
-It is good at systems with not too small noise and not too much unstableness, it does not require hyperbolicity.
-However, it can be expensive when the noise is small and unstableness is big.
+It works well when the noise level is not too small or the largest Lyapunov exponent is not too large; it also does not require hyperbolicity.
+The adjoint version solves a difficult version of variational data assimilation problem in chaotic systems with unknown parameters and partial observation, where the loss is a single long-time functional accounting for discrepancies in both the observations and the dynamics.
+However, it can be expensive when the noise is small and the largest Lyapunov exponent is large.
 Moreover, only part of the expression is derivative on the terminal densities.
 The path-kernel idea was also used in the Bismut-Elworthy-Li formula for the derivative with respect to the initial conditions [4, 9, 41].
-Our divergence-kernel method in [24] gives the score (derivative of marginal density) of random dynamical systems and SDEs.
-It works for non-hyperbolic and multiplicative noise systems.
-It is good at systems with not too much contraction and not too little noise.
+Our divergence-kernel method in [31] gives the score (spatial-derivative of marginal density) of random dynamical systems and SDEs.
+Compared to previous pathwise expressions for scores [45, 22, 26], the main advantage is that it works for multiplicative noise efficiently.
+It is good at systems with not too much contraction or not too little noise.
 However, it can be expensive when the noise is small and the contraction is strong.
-We use this result for the forward-backward diffusion models via nonlinear SDEs [18].
-This paper extends this result to linear responses.
+It fits into the conventional framework of score-based diffusion models, and it enables us to use nonlinear SDEs with multiplicative noise as the forward model, since it can compute the scores efficiently.
+This paper extends this result to linear responses, which further enables a new framework of generative models.
+1.2.2. Generative models
 Our new linear response formula opens up the possibilities of redesigning the current framework of diffusion models.
-The current structure of the diffusion model involves a forward process and a backward process [35, 16].
-Part of the reason is that previously we did not have a gradient tool to optimize marginal distributions with respect to the diffusion coefficient or under multiplicative noise.
-If we can directly compute such a gradient, we can directly optimize the diffusion process so that the marginal distribution matches the data.
-We will demonstrate this idea on some simple problems.
+At a high level, the bottleneck in ‘parametric SDE as a generative model’ is that the most direct training loss for i.i.d.
+data is the KL-divergence.
+But optimizing this loss requires differentiation of the model’s marginal density.
+For general SDEs this was typically intractable, where many classical tricks either break or become prohibitively high-variance.
+So previously we had to either replace the training loss or simplify the SDE.
+A first workaround is to keep using a parameterized SDE to learn a distribution, but replace KL-divergence training loss by the discrepancy over expectations of certain observables; this covers adversarial critics, kernel MMD, and moment matching [1, 17, 20].
+This new loss does not require evaluating the terminal marginal density or its parameter-gradient, so we can use the standard backpropagation to compute the gradient of this loss.
+However, in chaotic or unstable settings, we need to replace conventional backpropagation method by new methods such as the adjoint path-kernel method [29].
+More importantly, the fit of the data is limited by the chosen observable or the class holding the observable: the user needs to introduce an additional modeling choice (the observable/critic/kernel/features/weights) that mediates what aspects of the distribution are matched.
+A second workaround, prominent in diffusion models [44] and flow-matching methods [21, 49], is to introduce an auxiliary probability path: a designer-chosen time-indexed family of intermediate probability distributions $\{H_{t}\}_{t\in[0,T]}$ (typically $T=1$) connecting a simple base distribution $H_{0}$ to the data distribution $H_{T}$.
+The main object to be learned now is the score or a velocity field, and the corresponding training loss becomes the score/denoising loss or vector-field regression loss.
+Similar to the first workaround, this training loss can be differentiated by standard backpropagation.
+However, the auxiliary probability path is typically decided without consulting the prior structural knowledge about the model, so the prior knowledge does not automatically translate into computational savings.
+Even if we have a very simple SDE with a few parameters, the score is not necessarily simple, so the model for the score still needs a large network and potentially many training steps.
+A third workaround is to keep using the KL divergence as the loss, but restrict the model class so that the dependence of the distribution on the parameters becomes tractable.
+We give three examples.
+First, autoregressive models factorize the joint density as a product of conditional densities under a chosen ordering; this introduces an additional modeling choice (the ordering) and typically requires causal/masked architectural constraints to parameterize these high-dimensional conditionals [38].
+Second, in normalizing flows [7] one represents the model distribution through an invertible transport map, so that the change-of-variables formula expresses the density via a Jacobian determinant.
+Gradients of the resulting log-determinant terms are closely related to divergence-type expressions, connecting to the divergence (transfer-operator) method in linear response (see the review in Section˜1.2.1); in practice, moderately deep flows often require additional constraints or regularization to maintain good conditioning.
+Third, one may fix the diffusion coefficient, which makes likelihood-based gradients more tractable via likelihood-ratio/Girsanov-type estimators or related score-function identities.
+However, this narrows the model class: if the diffusion cannot be optimized, the induced SDE marginals may remain far from the data distribution.
+Summarizing, our divergence-kernel method enables the gradient tool to optimize KL divergence with respect to the diffusion coefficient or under multiplicative noise.
+With this, we can work with the simplest generative model, without introducing an auxiliary probability path or an auxiliary observable and without freezing the diffusion.
+This setting is common rather than exceptional: in mechanistic SDE modeling one often wants likelihood-style marginal fitting while allowing parameters in the diffusion; existing approaches typically avoid this combination.
 1.3. Structure of paper
@@
 Moreover, compared with our path-kernel methods, here we can compute the derivative of the density.
-Section˜5 proposes a diffusion model which involves only two forward processes.
-It uses the divergence-kernel formula (not the backpropagation method, which belongs to the path-perturbation method) to perform a gradient descent on the parameters of an SDE.
+Section˜5 proposes a parametric SDE generative model, DK-SDE (DK is short for divergence-kernel), which involves only two forward processes.
+It uses the divergence-kernel formula (not the backpropagation method, which belongs to the path-perturbation method) to perform a gradient descent on the parameters of an SDE model.
 Its loss function is the KL-divergence between the data and the marginal distribution of the SDE.
@@
 where $\left|g_{*}\right|$ is the determinant of the matrix.
+It is called a divergence due to an equivalent definition by contracting the Hessian (see [35]).
+Note that this is not the row-wise (or column-wise) divergence of a matrix.
 For discrete-time results in Section˜3.2, we assume that
@@
 The density of the initial distribution $h_{0}$, the probability kernel $k$, drift $f$, and diffusion $\sigma$ are $C^{1}$ jointly in $\gamma$ and $x$.
+The diffusion field $\sigma^{\gamma}(x)$ is uniformly bounded away from zero.
 Our derivation is rigorous for the discrete-time case, but is formal when passing to the continuous-time limit, so we do not list the technical assumptions here, and leave the rigorous proof to a later paper.
 We just assume that all integrations, averages, and change of limits are legit.
-We still organize our continuous-time results by lemmas and theorems, but this is just a style to organize the paper and does not imply rigorousness.
+All numerical experiments use the time-discretized formulas; the continuous-time statements serve as guiding limits.
 3. Derivation of the divergence-kernel formula for linear responses
@@
 Another scenario where this approach might be more useful is that the noise is low-dimensional.
-Bot situations incur other difficulties, so we will defer to later papers.
+Both situations incur other difficulties, so we will defer to later papers.
 3.2. Discrete-time divergence-kernel formula for linear responses
@@
 [N-step forward divergence-kernel formula for score [31]] For any backward adapted process $\{\alpha_{n}\}_{n=0}^{N}$, let $\{\nu_{n}\}_{n=0}^{N}$ be the forward covector process
-$\begin{split}\nu_{0}=\nabla\log h_{0}(x_{0}),\quad\textnormal{}\quad\nu_{n+1}=(1-\alpha_{n+1})g_{b_{n}}^{*-1}\left(\nu_{n}-\textnormal{div}g_{b_{n}*}(x_{n})\right)+\frac{\alpha_{n+1}\nabla\log k(b_{n})}{\sigma_{n}(x_{n})}.\end{split}$
+$\begin{split}\nu_{0}=\nabla\log h_{0}(x_{0}),\qquad\nu_{n+1}=(1-\alpha_{n+1})g_{b_{n}}^{*-1}\left(\nu_{n}-\textnormal{div}g_{b_{n}*}(x_{n})\right)+\frac{\alpha_{n+1}\nabla\log k(b_{n})}{\sigma_{n}(x_{n})}.\end{split}$
 Then for any $N\geq 0$,
@@
 Repeat this on $\delta\log h^{\gamma}_{n}(x_{n})$ for each $n$ to prove the lemma. ∎
-Note that we need the score of $h_{0}$, or $\nu_{0}$. Only for large $T$ and systems with fast decorrelation, we can use any $\nu_{0}$ and the impact would be negligible. See ˜Remark for more details.
+Note that we need the score of $h_{0}$, or $\nu_{0}$.
+Only for large $T$ and systems with fast decorrelation, we can use any $\nu_{0}$ and the impact would be negligible.
+See Section˜3.4 for more details.
 3.3. Preparations for SDEs
@@
 $\begin{split}\left[\nabla F\right]_{ij}=\frac{\partial F^{i}}{\partial x^{j}}=:\partial_{j}F^{i},\quad\quad\left[{\Delta B}\nabla\sigma^{T}\right]_{ij}=\frac{\partial\sigma}{\partial x^{j}}{\Delta B}^{i}=:\partial_{j}\sigma{\Delta B}^{i}.\end{split}$
+We keep only terms of order $\Delta t$} in expectation and drop higher-order terms since they integrate to zero; this is standard in Ito calculus.
 Ignoring terms that go to zero in the continuous-time limit, [31] derived explicit expressions of some terms for time-discretized SDEs,
@@
 Then we recall the expression of the score function at $\gamma=0$.
-{lemma}
-[formal divergence-kernel fomula for score of SDEs [24]] For the SDE in $\mathbb{R}^{M}$,
+[…]
+For the SDE in $\mathbb{R}^{M}$,
 $dx_{t}=F_{t}(x_{t})dt+\sigma_{t}(x_{t})dB_{t},$
@@
 First, we give an expression of $\delta\log|g^{\gamma}_{b_{n}*}|$, then we derive the linear response formulas.
-{lemma}
+[…]
 Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$,
 $\begin{split}\delta\log|g^{\gamma}_{b_{n}*}|\approx\textnormal{div}\delta F_{n}^{\gamma}{\Delta t}+\nabla\delta\sigma_{n}^{\gamma}\cdot{\Delta B}_{n}-\nabla\sigma_{n}^{T}\nabla\delta\sigma_{n}^{\gamma}{\Delta t}.\end{split}$
-Proof.
+Derivation.
 Substitute the expression in Equation˜8, we get
@@
 $\begin{split}(\nabla\delta\sigma^{\gamma}\cdot{\Delta B})(\nabla\sigma\cdot{\Delta B})=\sum_{i}\sum_{j}\partial_{i}\delta\sigma^{\gamma}{\Delta B}^{i}\partial_{j}\sigma{\Delta B}^{j}\approx\sum_{i}\partial_{i}\delta\sigma^{\gamma}\partial_{i}\sigma{\Delta t}=\nabla\sigma^{T}\nabla\delta\sigma^{\gamma}{\Delta t}.\end{split}$
-Then substitute to prove the corollary. ∎
-{lemma}
-[expression of $\delta x^{\gamma}_{n}$] Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$,
+Then substitute to derive the formula. ∎
+[…]
+Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$,
 $\begin{split}\delta x^{\gamma}_{n}=-g^{-1}_{b_{n}*}\delta g^{\gamma}_{b_{n}}(x_{n})\approx-\delta F^{\gamma}_{n}{\Delta t}-\delta\sigma^{\gamma}_{n}{\Delta B}_{n}+\delta\sigma^{\gamma}_{n}\nabla\sigma_{n}{\Delta t}.\end{split}$
@@
 For a discrete-time system, this term would involve solving a $M$-dimensional linear equation system.
-Proof.
+Derivation.
 By Equation˜7,
@@
 Then substitute to prove the lemma. ∎
-{lemma}
+[…]
 Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$, let $\Delta\sigma$ denote the Laplacian of $\sigma$, then
 $\begin{split}\nabla\log\left|g_{b_{n}*}\right|\cdot\delta x^{\gamma}_{n}\approx-\delta\sigma^{\gamma}_{n}\Delta\sigma_{n}{\Delta t}.\end{split}$
-Proof.
+Derivation.
 By Equation˜8 and Section˜3.4,
@@
 Then substitute to prove the lemma. ∎
-With these, we can formally pass Definition˜1 to the continuous-time limit.
-See 1.1
-Proof.
-Substitute Sections˜3.4, 3.4 and 3.4 into Definition˜1, we get
+With these, we can formally pass Definition˜1 to the continuous-time limit to derive the main formula of this paper.
+Derivation.
+(of Section˜1.1) Substitute Sections˜3.4, 3.4 and 3.4 into Definition˜1, we get
 (10) $\begin{split}\delta\log h^{\gamma}_{N}(x_{N})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}+\sum_{n=0}^{N-1}\nu_{n}\cdot\delta x^{\gamma}_{n}-\nabla\log|g_{b_{n}*}|\cdot\delta x^{\gamma}_{n}-\delta\log|g^{\gamma}_{b_{n}*}|\middle|x_{N}\right]\\ \approx\mathbb{E}\left[\delta\log h^{\gamma}_{0}+\sum_{n=0}^{N-1}\nu_{n}\cdot(-\delta F^{\gamma}_{n}{\Delta t}-\delta\sigma^{\gamma}_{n}{\Delta B}_{n}+\delta\sigma^{\gamma}_{n}\nabla\sigma_{n}{\Delta t})\right.\\ \left.+\delta\sigma^{\gamma}_{n}\Delta\sigma_{n}{\Delta t}-\textnormal{div}\delta F_{n}^{\gamma}{\Delta t}-\nabla\delta\sigma_{n}^{\gamma}\cdot{\Delta B}_{n}+\nabla\sigma_{n}^{T}\nabla\delta\sigma_{n}^{\gamma}{\Delta t}\middle|x_{N}\right].\end{split}$
@@
 We need to further assume that $F_{t}$ and $\sigma_{t}$ are stationary, and that there is $T^{\prime}$ such that $\alpha_{t}$ only depends on $x_{t}\sim x_{T^{\prime}}$, so $(\alpha,\nu,x)$ are jointly stationary.
-{corollary} [formal ergodic divergence-kernel linear response] For any $C^{2}$ observable function $\Phi:\mathbb{R}^{M}\rightarrow R$, let $h^{\gamma}$ be the stationary measure under parameter $\gamma$, then the linear response of the averaged observable can be expressed almost surely on one path:
+[…]
+For any $C^{2}$ observable function $\Phi:\mathbb{R}^{M}\rightarrow R$, let $h^{\gamma}$ be the stationary measure under parameter $\gamma$, then the linear response of the averaged observable can be expressed almost surely on one path:
 […]
@@
 We may need to use all three methods (path, kernel, and divergence) together to obtain an ultimate solution, as proposed in [32].
+The divergence-kernel method contains second derivatives.
+To speed up our code, note that we do not really need the matrices such as $\nabla^{2}\sigma$ and $\nabla F^{T}$, we only need their products with certain vectors.
+Such matrix-vector products can be implemented much more efficiently in high dimensions (such as the jvp function in JAX) than forming the matrix, especially with sparse structures.
 Compared to the path-kernel method in [30, 29], the divergence-kernel method has the following advantages:
@@
     For dynamical systems generated from discretized PDEs, typically $|\lambda_{min}|>|\lambda_{max}|$, where $\lambda_{max}$ and $\lambda_{min}$ are the largest and the smallest (most negative) Lyapunov exponents; hence, the path-kernel method requires a smaller $\alpha$, so the estimator has smaller size and the average converges faster.
-4. Numerical examples
+4. Numerical examples for linear responses
 4.1. 1-dimensional process with multiplicative noise
 We use Section˜1.1 to compute the linear response of the one-dimensional process with multiplicative noise. The SDE is
-$\begin{split}dx=\beta_{t}(\gamma-x)dt+\sqrt{\beta_{t}}(0.5+\exp\left((\gamma-x)^{2}\right)dB,\quad\textnormal{where}\quad\beta_{t}:=1+3t.\end{split}$
-The initial distribution is $x_{0}\sim N(0,1)$. Set $T=0.1$, ${\Delta t}=0.01$, $\alpha=10$. Note that the problem is symmetric for $\gamma$.
-We are only interested in terminal conditions $x_{T}$ in the interval [2,2][-2,2], which is further partitioned into $10$ subintervals.
+$\begin{split}dx=\beta_{t}(\gamma-x)dt+\sqrt{\beta_{t}}\left(0.5+\exp\left((\gamma-x)^{2}\right)\right)dB,\quad\textnormal{where}\quad\beta_{t}:=1+3t.\end{split}$
+The initial distribution is $x_{0}\sim N(0,1)$. Set $T=1$, ${\Delta t}=0.01$, $\alpha=10$. Note that the problem is symmetric for $\gamma$.
+We are only interested in terminal states $x_{T}$ in the interval [2,2][-2,2], which is further partitioned into $10$ subintervals.
 We simulate many sample paths, then count the number of paths whose $x_{T}$ fall into a specific subinterval, to obtain the empirical distribution of $x_{T}$ on subintervals.
@@
 Each subinterval contains at least 180 samples.
-The $\delta\log h_{T}$ and $\nabla\log h_{T}$ computed for different $\gamma$ and $x$’s are shown in Figure˜1.
+The $\delta\log h_{T}$ and $\nabla\log h_{T}$ computed for different $\gamma$ and $x$ are shown in Figure˜1.
 Both correctly reflect the trend of $\log h_{T}$.
 Figure 1.
-Divergence-kernel method for the linear responses and scores of the 1-dimensional SDE with multiplicative noise, $T=0.3$.
+Divergence-kernel method for the linear responses and scores of the 1-dimensional SDE with multiplicative noise, $T=1$.
 The dots are $\log h_{T}$.
-Each short line is a linear response or score computed by the divergence-kernel algorithm, averaged on paths whose terminal conditions $x_{T}$ fall into the same subinterval.
+Each short line is a linear response or score computed by the divergence-kernel algorithm, averaged on paths whose terminal states $x_{T}$ fall into the same subinterval.
 Left: linear responses $\delta\log h_{T}$.
@@
 4.2. 40-dimensional Lorenz 96 system with multiplicative noise
-We use ˜Remark to compute the linear response of the stationary density of the Lorenz 96 model [14]. The dimension of the system is $M=40$. The SDE is
-$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+8+\gamma-0.01(x^{i})^{2}\right)dt+0.5+\exp\left((x-\gamma)^{2}\right)dB^{i},$
-$\textnormal{where}\quad i=1,\ldots,M;\quad x_{0}\sim{\mathcal{N}}(0,I).$
-Here $i$ labels different directions in $\mathbb{R}^{M}$, and it is assumed that $x^{-1}=x^{M-1},x^{0}=x^{M}$ and $x^{M+1}=x^{1}$.
+We use Section˜3.4 to compute the linear response of the stationary density of the Lorenz 96 model [23]. The dimension of the system is $M=40$. The SDE is
+$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+8+\gamma-0.01(x^{i})^{2}\right)dt+\left(0.5+\exp\left(|x-\gamma|^{2}\right)\right)\,dB^{i},$
+$\textnormal{where}\quad i=0,\ldots,M-1;\quad x_{0}\sim{\mathcal{N}}(0,I).$
+Here $i$ labels different directions in $\mathbb{R}^{M}$, and We use cyclic indexing modulo $M$.
 We added noise and the $0.01(x^{i})^{2}$ term, which prevents the noise from carrying us to infinitely far away.
@@
 A typical orbit is in Figure˜2.
-Then we use the ergodic linear response formula ˜Remark to compute the linear response of the observable averaged on the stationary measure.
+Then we use the ergodic linear response formula Section˜3.4 to compute the linear response of the observable averaged on the stationary measure.
 The averaged observable $\Phi_{avg}$ and the linear response $\delta\int\Phi h^{\gamma}$ are averaged over 7 orbits of length $T=400$.
@@
 Figure 2. Left: plot of $x^{0}_{t},x^{1}_{t}$ from a typical orbit of time length $1.5$. Right: linear response computed by divergence-kernel method.
-5. A simple forward-only diffusion model
-We consider the problem in which we are given $N_{data}$ many samples $y_{k}\in\mathbb{R}^{M}$, whose true distribution is unknown.
-Then we use gradient descent to find an SDE such that it maps the Gaussian distribution to match the data.
-First, we describe how to generate samples. Note that the information in this paragraph will not be available when we infer the SDE. The samples are generated by the SDE
+5. Parametric SDE generative model trained by the divergence-kernel method (DK-SDE)
+We consider the problem in which we are given $N_{\mathrm{data}}$ many samples $y_{k}\in\mathbb{R}^{M}$, whose true distribution is unknown.
+Then we use the divergence-kernel method to compute the gradient, and use gradient descent to train a parametric generative SDE model which maps the Gaussian distribution to match the data.
+5.1. Problem and algorithm description
+We describe the problem and our model, then we define loss function and explain how to compute its gradient, and how to update the parameters by gradient descent.
+First, we describe how to generate samples. Note that the information in this paragraph will not be available when we train the model. The samples are generated by the SDE
 $\begin{split}dx_{t}=F^{\gamma_{true}}_{t}(x_{t})dt+\sigma^{\gamma_{true}}_{t}(x_{t})dB_{t},\end{split}$
-where the true parameters $\gamma_{true}:=[\gamma_{true}^{0},\ldots,\gamma_{true}^{N_{\gamma}}]$.
+where the true parameters $\gamma_{true}:=[\gamma_{true}^{0},\ldots,\gamma_{true}^{N_{\gamma}-1}]$.
 The samples are $y:=x_{T}$ obtained at the end of independent paths.
-Denote the empirical measure, which we try to match, as
-$\begin{split}P(\cdot):=\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\delta_{y_{k}}(\cdot).\end{split}$
-Our framework can work on more general data, but here we restrict ourselves to the simple case, so that after optimization, we can check our result with the true parameters.
-We want to fit several unknown parameters $\gamma=[\gamma^{0},\ldots,\gamma^{N_{\gamma}}]$ in an SDE such that its marginal distribution matches the data. The prototype of the SDE is
+Denote the target distribution (the one we try to match) as $P$; when the target is empirical distribution prescribed by a set of data, $P$ is a sum of Dirac delta functions.
+Our framework can work on more general data not necessarily generated by an SDE, but here we restrict ourselves to this simple case, so that after optimization, we can check our result with the true parameters.
+We want to fit several unknown parameters $\gamma=[\gamma^{0},\ldots,\gamma^{N_{\gamma}-1}]$ in the generative SDE model such that its marginal distribution matches the data.
+The prototype of the SDE is
 $\begin{split}dx_{t}=F^{\gamma}_{t}(x_{t})dt+\sigma^{\gamma}_{t}(x_{t})dB_{t}.\end{split}$
 Denote the density at $T$ by $h^{\gamma}$ and measure by $H^{\gamma}$ – this is the distribution that we want to optimize.
+In the current paper, we use simple models, such as one based on the Lorenz 96 system.
+These simple models can be interpreted as the result of the prior knowledge we had about how the data was generated.
+Such prior knowledge can be naturally incorporated into the SDE model.
+In the future, we will test on problems with less prior knowledge, so we will need more flexible models to express $F$ and $\sigma$, such as neural networks.
+Also, note that the gradient used in the optimization is not computed by backpropagation; rather, it only involves a covector process that runs forward in time and uses much less memory.
+We also remark that we can choose to optimize the stationary distribution using Section˜3.4, which may be more efficient for some cases.
+5.2. KL divergence as loss and gradient descent
+Our training loss is the KL divergence, but we never need to evaluate the KL-divergence $D_{\mathrm{KL}}(P\|H^{\gamma})$; we only need its parameter-gradient.
+When $P$ admits a density $p$ (w.r.t.
+Lebesgue measure), the KL-divergence is
+$D_{\mathrm{KL}}(P\|H^{\gamma})=\int p(y)\log\frac{p(y)}{h^{\gamma}(y)}\,dy=\underbrace{\int p(y)\log p(y)\,dy}_{\text{independent of }\gamma}-\int p(y)\log h^{\gamma}(y)\,dy.$
+Hence, its parameter-derivative is
+$\delta D_{\mathrm{KL}}(P\|H^{\gamma})=-\int_{\mathbb{R}^{M}}p(y)\delta\log h^{\gamma}(y)\,dy.$
+If $P$ is given only through samples $\{y_{k}\}_{k=1}^{N_{\mathrm{data}}}\sim P$ (or is strictly singular), we optimize the same $\gamma$-dependent term (cross-entropy), namely
+$\mathcal{L}(\gamma):=-\frac{1}{N_{\mathrm{data}}}\sum_{k=1}^{N_{\mathrm{data}}}\log h^{\gamma}(y_{k}),\qquad\delta\mathcal{L}(\gamma)=-\frac{1}{N_{\mathrm{data}}}\sum_{k=1}^{N_{\mathrm{data}}}\delta\log h^{\gamma}(y_{k}).$
+We can compute the cross-entropy, but there is little benefit in doing so here, and it would require a few additional steps.
+In this paper, we judge the accuracy of the trained model by the discrepancy between the trained parameters and the ground-truth parameters; this criterion is complementary to the KL/cross-entropy.
+Then we need to decide how to compute $\delta\log h^{\gamma}$ at the prescribed locations $y_{k}$.
+It is difficult to obtain paths ending exactly at $y_{k}$, so we need an approximation.
 For each $\gamma$ during optimization, we simulate $L$-many paths by this SDE, and we label each path by $l$, in particular, the terminal location is denoted by $x_{T,l}$.
-In future research, we can seek to use more flexible prototypes, such as neural networks.
-However, note that the gradient used in the optimization is not computed by backpropagation; rather, it only involves a covector process that runs forward in time and costs much less memory.
-We also remark that we can choose to optimize the stationary distribution using ˜Remark, which may be more efficient for some cases.
-The difference between the generated distribution and the data is measured by the KL-divergence
-$\begin{split}D_{\text{KL}}(P\parallel H^{\gamma}):=\int_{y\in{\mathbb{R}^{M}}}\log{\frac{P(dy)}{H^{\gamma}(dy)}}\,P(dy).\end{split}$
-Its parameter-derivative is
-$\begin{split}\delta D_{\text{KL}}(P\parallel H^{\gamma})=-\int_{y\in{\mathbb{R}^{M}}}\frac{\delta H^{\gamma}(dy)}{H^{\gamma}(dy)}\,P(dy)=-\int_{y\in{\mathbb{R}^{M}}}\frac{\delta h^{\gamma}}{h^{\gamma}}(y)\,P(dy)\\ =-\int_{y\in{\mathbb{R}^{M}}}\delta\log h^{\gamma}(y)\,P(dy)=-\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\delta\log h^{\gamma}(y_{k}).\end{split}$
-With the new formula in Section˜1.1, we have an expression of $\delta\log h^{\gamma}$ at $x_{T,l}$’s, which enables the gradient optimization on this KL divergence.
-Then we need to decide how to compute $\delta\log h^{\gamma}$ at the prescribed locations $y_{k}$’s.
-It is difficult to obtain paths ending exactly at $y_{k}$, so we need an approximation.
-We approximate $y_{k}$ by its $N_{neighbor}$-nearest neighbors in $\{x_{T,l}\}_{l=1}^{L}$, and therefore $\delta\log h^{\gamma}(y_{k})$ is approximated by the average of the divergence-kernel expression in Section˜1.1 over those $N_{neighbor}$ paths.
+We approximate $y_{k}$ by its $N_{\mathrm{neighbor}}$-nearest neighbors in $\{x_{T,l}\}_{l=1}^{L}$, and therefore $\delta\log h^{\gamma}(y_{k})$ is approximated by the average of the divergence-kernel expression in Section˜1.1 over those $N_{\mathrm{neighbor}}$ paths.
 Hence,
-$\begin{split}\delta D_{\text{KL}}(P\parallel H^{\gamma})\approx\hat{\delta}D_{\text{KL}}(P\parallel H^{\gamma}):=-\frac{1}{N_{data}}\sum_{k=1}^{N_{data}}\frac{1}{N_{neighbor}}\sum_{x_{T,l}\in N(y_{k})}\mathcal{T}(x_{T,l}),\end{split}$
+(11) $\begin{split}\delta D_{\mathrm{KL}}(P\|H^{\gamma})\approx\hat{\delta}D_{\mathrm{KL}}(P\|H^{\gamma}):=-\frac{1}{N_{\mathrm{data}}}\sum_{k=1}^{N_{\mathrm{data}}}\frac{1}{N_{\mathrm{neighbor}}}\sum_{x_{T,l}\in N(y_{k})}\mathcal{T}(x_{T,l}),\end{split}$
 where $N(y_{k})$ is the neighbors of $y_{k}$, and $\mathcal{T}(x_{T,l})$ is the expression in Section˜1.1 on the path ending at $x_{T,l}$.
 After approximately computing the parameter-derivative, we update the parameter by gradient descent,
-$\begin{split}\gamma_{n+1}=\gamma_{n}-\eta\hat{\delta}D_{\text{KL}}(P\parallel H^{\gamma}).\end{split}$
+$\begin{split}\gamma_{n+1}=\gamma_{n}-\eta\hat{\delta}D_{\mathrm{KL}}(P\|H^{\gamma}).\end{split}$
 For simplicity, we keep $\eta$ constant, but we can also change it after each update of $\gamma$.
-5.1. A 1-d problem with 2 parameters
-First, we try a 1-dimensional problem. The SDE is
-$\begin{split}dx^{i}=\left(\gamma^{0}-x\right)dt+\left(0.5+\exp\left(-(x-\gamma^{1})^{2}\right)\right)dB,\quad\textnormal{where}\quad x_{0}\sim{\mathcal{N}}(0,1).\end{split}$
+We can give a direct interpretation of the overall optimization strategy.
+Before each update, we simulate $L$ many sample paths, then we decide which of these paths’ terminal state is closest to the data (this is the effect of choosing $N_{\mathrm{neighbor}}$-nearest neighbors).
+On the other hand, the divergence-kernel linear response formula tells us how changing parameters affects the marginal log-densities at those terminal states.
+Combining both facts, we can now update parameters in the direction which increases the marginal log-densities on those terminal states close to the data.
+Hence, overall, each parameter update moves the marginal distribution of the SDE towards the data.
+This interpretation also suggests that the best choice of $N_{\mathrm{neighbor}}$ is often $1$ or a small number.
+Indeed, increasing $N_{\mathrm{neighbor}}$ replaces the ideal conditional expectation $\mathbb{E}[\cdot\mid x_{T}=y_{k}]$ by a local average over nearby terminal states.
+While this can reduce Monte-Carlo variance, it also introduces a bias: points in high-density regions of the sample cloud tend to be selected as neighbors more frequently than points near the periphery.
+As a result, when $N_{\mathrm{neighbor}}$ is too large, the update may overweight central samples and encourage an artificial contraction of the learned distribution toward its center, even when the parameters are already close to optimal.
+For example, in the idealized situation where the simulated terminals match the data exactly, taking $N_{\mathrm{neighbor}}=1$ yields an unbiased 1-to-1 neighbor assignment, whereas $N_{\mathrm{neighbor}}>1$ mixes in off-target terminals and reintroduces geometric bias toward locally denser regions.
+Therefore we typically take $N_{\mathrm{neighbor}}=1$ (or a small value) unless additional smoothing is needed for stability.
+We stress that the nearest-neighbor step is merely a simple numerical device for approximating the conditional expectation appearing in the divergence-kernel formula.
+One could replace it by other local averaging schemes without changing the underlying update principle.
+Therefore, the more substantive question is not the optimality of nearest-neighbor, but whether the resulting update direction admits the correct interpretation: namely, that it increases the model marginal log-density in regions supported by the data.
+5.3. A 1D problem with 2 parameters
+First, we try a 1-dimensional problem. The SDE model is
+$\begin{split}dx=\left(\gamma^{0}-x\right)dt+\left(0.5+\exp\left(-(x-\gamma^{1})^{2}\right)\right)dB,\quad\textnormal{where}\quad x_{0}\sim{\mathcal{N}}(0,1).\end{split}$
 The true parameter used to generate the data is $\gamma_{true}=[0,0]$.
 The initial parameter in the optimization is $\gamma_{0}=[5,1]$.
-The number of data points is $N_{data}=200$.
+The number of data points is $N_{\mathrm{data}}=200$.
 The number of samples generated for each update of $\gamma$ is $L=200$.
 The stepsize (or learning rate) is $\eta=1$.
-The number of neighbors used to approximate each $y_{k}$ is $N_{neighbor}=5$.
+The number of neighbors used to approximate each $y_{k}$ is $N_{\mathrm{neighbor}}=5$; we will test the effect of $N_{\mathrm{neighbor}}$ in the next example with higher dimension $M$.
+Now we decide $\alpha$.
+For simplicity, we let it be a constant.
+To decide its value, recall that $\alpha$ is used to temper the growth of the covector process $\nu$.
+In view of Section˜1.1, the homogeneous part of $\nu$ grows like
+$\begin{split}d\nu=(\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha)\nu dt+\cdots\end{split}$
+If $\alpha$ is too small then $\nu$ would grow exponentially fast; we do not want this.
+For this 1D problem, we can directly compute these terms.
+We want the homogeneous part of $\nu$ to decay, which can be achieved if
+$\begin{split}\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}=4\exp\left(-2(x-\gamma^{1})^{2}\right)(x-\gamma^{1})^{2}+1\leq 2/e+1\leq 3\leq\alpha\end{split}$
+So we can just set $\alpha$ to be somewhat larger than $3$, for example, we set $\alpha=5$.
+For more complicated situations, we can just try a few $\alpha$ and pick the one whose corresponding $\nu$ is moderate.
 We start from $\gamma_{0}$ and use gradient descent to minimize the KL-divergence.
-The histograms of $y_{k}$’s and $x_{T,l}$’s for $\gamma_{0}$ and $\gamma_{10}$, the parameter after 10 updates, are shown in Figure˜3.
+The histograms of data $\{y_{k}\}$ and samples $\{x_{T,l}\}$ for $\gamma_{0}$ and $\gamma_{10}$, the parameter after 10 updates, are shown in Figure˜3.
 As we can see, the two distributions get much closer during the optimization.
 More specifically, the error $\gamma-\gamma_{true}$ is reduced from $[5,1]$ to [0.05,0.09][-0.05,0.09].
-The computational time for all 10 updates on a 4-core CPU is about 2 minutes, but note that our implementation is far from optimal.
-Figure 3. Histograms of $y_{k}$’s and $x_{T,l}$’s. Left: plot at $\gamma_{0}$. Right: at $\gamma_{10}$.
-5.2. A 5-d problem with 6 parameters
-We try a 5-dimensional Lorenz 96 system. The SDE is
-$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+\gamma^{i}-0.01(x^{i})^{2}\right)dt+\left(0.5+\exp\left((x-\gamma^{M}\mathbbm{1})^{2}\right)\right)dB^{i},$
+Figure 3. Histograms of data $\{y_{k}\}$ and samples $\{x_{T,l}\}$. Left to right: plots at $\gamma_{0}$, $\gamma_{2}$, and $\gamma_{10}$.
+5.4. A 5D problem with 6 parameters
+We try a 5-dimensional Lorenz 96 system. The SDE model is
+$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+\gamma^{i}-0.01(x^{i})^{2}\right)dt+\left(0.5+\exp\left(-(x-\gamma^{M}\mathbbm{1})^{2}\right)\right)dB^{i},$
 $\textnormal{where}\quad i=0,\ldots,M-1;\quad M=5;\quad\mathbbm{1}=[1,1,1,1,1];\quad x_{0}\sim{\mathcal{N}}(0,I).$
@@
 The initial parameter in the optimization is $\gamma_{0}=[0,0,0,0,0,0]$.
-The number of data points is $N_{data}=200$.
+The number of data points is $N_{\mathrm{data}}=200$.
 The number of samples generated for each update of $\gamma$ is $L=200$.
 The stepsize (or learning rate) is $\eta=1$.
-The number of neighbors used to approximate each $y_{k}$ is $N_{neighbor}=5$.
+The number of neighbors used to approximate each $y_{k}$ is $N_{\mathrm{neighbor}}=5$.
 We start from $\gamma_{0}$ and use gradient descent to minimize the KL-divergence.
-The history of the $l^{2}$-distance $|\gamma_{n}-\gamma_{true}|$ for $n=0$ to $50$ is shown in Figure˜4.
+The history of $|\gamma_{n}-\gamma_{true}|^{2}/N_{\gamma}$ for $n=0$ to $50$ is shown in Figure˜4.
 As we can see, $\gamma_{n}$ gets closer to $\gamma_{true}$ during optimization, with $\gamma_{48}$ achieving the minimum distance.
@@
 $\begin{split}\gamma_{0}-\gamma_{true}=[-5,-6,-7,-8,-9,-2];\\ \gamma_{48}-\gamma_{true}=[-0.33,0.40,-0.26,-0.20,-0.24,-0.36].\end{split}$
-The computational time for all 50 updates on a 4-core CPU is about 2 minutes, but note that our implementation is far from optimal.
-Figure 4. History of $|\gamma_{n}-\gamma_{true}|$ for the 5-d diffusion model.
+Figure 4. History of $|\gamma_{n}-\gamma_{true}|^{2}/N_{\gamma}$ for the 5D DK-SDE model.
+5.5. A 20D problem with 21 parameters
+We try a 20-dimensional system.
+More importantly, we run some convergence test against several hyper-parameters and record computer time.
+The SDE model is similar to the one in Section˜5.4, except for that now we set dimension $M=20$ and $N_{\gamma}=21$.
+We set $\gamma_{\mathrm{true}}\in\mathbb{R}^{N}_{\gamma}$ to ramp linearly from $5$ to $9$ over the first ten components and then ramp linearly back from $9$ to $5$ over the next ten components; additionally, we set the last entry $\gamma_{\mathrm{true},\,N_{\gamma}-1}=1$.
+The initial guess of all parameters are still zeros.
+The number of data points is $N_{\mathrm{data}}=200$.
+We vary several hyperparameters; when sweeping one parameter, we keep the others fixed at their default values.
+In Figure˜5 we plot the convergence history of $|\gamma_{n}-\gamma_{\mathrm{true}}|^{2}/N_{\gamma}$ for $n=0,\dots,1000$ under different choices of hyperparameters.
+Specifically, we sweep $\eta$, $N_{\mathrm{neighbor}}$, and $L$, while holding the remaining two at the defaults $\eta=0.1$, $N_{\mathrm{neighbor}}=1$, and $L=3000$.
+Smaller $\eta$ leads to slower convergence but a better final error; in future work, it may therefore be beneficial to decrease $\eta$ according to a schedule.
+Using $N_{\mathrm{neighbor}}=1$ introduces some oscillations but yields the best overall result, which is encouraging since this choice does not depend on the dimension $M$.
+Finally, increasing the sample size $L$ improves the results, as expected.
+Refer to captionRefer to captionRefer to caption
+Figure 5. Convergence history for the 20D DK-SDE model. From left to right: sweep over learning rate $\eta$, $N_{\mathrm{neighbor}}$, and number of samples $L$.
+Then we plot the data and generated samples under the default hyperparameters in Figure˜6.
+The wall time for 1000 updates on an RTX 5060 GPU is about 201 seconds.
+As intended, the sample distribution moves closer to the data distribution during optimization.
+More specifically, the scatter plots indicate good agreement at least in the first and second moments, which is a natural consequence of distributions getting closer.
+Refer to captionRefer to captionRefer to caption
+Figure 6.
+Scatter plot of $(x^{0},x^{1})$ for samples generated by the DK-SDE model and the given data.
+From left to right: samples generated by SDE with initial parameter, after 70 updates, after 500 updates.
+Direct baselines are unavailable for this problem.
+Methods based on differentiating through trajectories (adjoint/tangent backpropagation) are numerically ineffective once the dynamics becomes unstable/nonhyperbolic (often already at moderate horizons in our setting), while likelihood-ratio constructions typically do not apply to parameters in the diffusion or suffer prohibitive variance.
+Other approaches (e.g.
+discrepancy or feature matching) optimize different objectives.
 Declarations
@@
 The second term in Section˜A.1 can be very large for high-dimensional noise since it involves $M$.
-For now, we still use noise in all directions, so we try to avoid using the kernel formula for parameter derivative.
+For now, we still use noise in all directions, so we try to avoid using the kernel formula for parameter-derivative.
 The kernel formula for the spatial derivative does not have terms involving $M$ [31], so we can still use that.
@@
 [likelihood ratio formula for linear response of SDEs] If $\delta\sigma\equiv 0$,
-$\begin{split}\nabla\log h_{T}(x_{T})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}(x_{0})+\int_{t=0}^{T}\frac{1}{\sigma_{t}(x_{t})}\delta F^{\gamma}(x_{n})\cdot dB\middle|x_{N}\right].\end{split}$
+$\begin{split}\delta\log h_{T}(x_{T})=\mathbb{E}\left[\delta\log h^{\gamma}_{0}(x_{0})+\int_{t=0}^{T}\frac{1}{\sigma_{t}(x_{t})}\delta F^{\gamma}(x_{t})\cdot dB\middle|x_{T}\right].\end{split}$